### v1

# itemizestditemize

Abstract
We present generative predictive control (GPC), a learning control framework that (i) clones a generative diffusion-based policy from expert demonstrations, (ii) trains a predictive action-conditioned world model from both expert demonstrations and random explorations, and (iii) synthesizes an online planner that ranks and optimizes the action proposals from (i) by looking ahead into the future using the world model from (ii). Crucially, we show that conditional video diffusion allows learning (near) physics-accurate visual world models and enable robust visual foresight. Focusing on planar pushing with rich contact and collision, we show GPC dominates behavior cloning across state-based and vision-based, simulated and real-world experiments.
Figure 1: Generative Predictive Control (GPC). Given visual (or state) observations, a generative policy—trained from behavior cloning—proposes an action chunk. Conditioned on the action chunk and past observations, a predictive world model—trained from diverse trajectories—predicts the future observations in “imagination”. In (a) GPC-RANK, the policy outputs multiple action proposals and for each proposal the world model predicts its future and associated reward. The proposal with highest reward in executed. In (b) GPC-OPT, the policy outputs a single action proposal and multiple gradient ascent steps are performed to optimize the action proposal for higher reward. Both GPC-RANK and GPC-OPT dominate pure behavior cloning.
Equal contribution
I Introduction
State-of-the-art robot control policies are synthesized using two major paradigms:
  • Predictive control. With access to a predictive dynamical model of the robot and its physical environment, optimal control (OC) and (deep) reinforcement learning (RL) are the workhorses for designing performant controllers in locomotion [57], manipulation [4], and aviation [1, 31]. While OC typically designs controllers based on analytic dynamical models and online optimization [30], RL can learn a policy via interactions with a complicated and black-box dynamical model [49], either through simulation or in the real world. The nature of predictive control is to look ahead, by anticipating how the environment will change and how much reward will be collected.
  • Generative control. Motivated by the success of generative modeling in vision and language, generative control—learning a controller from expert demonstrations using behavior cloning with generative models—is rapidly rising to prominence [15, 52]. Due to the abundance of demonstrations and the unprecedented distribution learning capacity of generative models, generative control has led to robot foundation policies capable of solving diverse tasks from vision and language [21]. The nature of generative control is to look back, making decisions based on expert policies in the past.
Each paradigm has its advantages and limitations. When dynamics modeling is accurate and tractable, predictive control delivers precision and robustness (e.g., Boston Dynamics parkour and SpaceX rocket landing). However, in many cases (e.g., manipulation with contact [35] and deformable objects [65], vision-based control [48]), the dynamics can be inaccurate or computationally expensive for optimization (as in OC) and interaction (as in RL). Generative control, on the other hand, circumvents the need for careful dynamics modeling and directly learns the policy via generative modeling of the expert demonstrations. However, generative control suffers from distribution shifts (due to its supervised learning nature [38, 21]) and lack of precision and robustness (i.e., they can do lots of tasks but do few of them as well as specialized controllers [19, 10]). Moreover, whether the “scaling law” can address these issues is questionable [7, 63, 62].
Humans, nevertheless, look back and look ahead when making decisions. We look back at similar past experiences to shortlist a set of proper actions, and look ahead—unroll the environment dynamics in “imagination”—to predict which action plan is optimal in terms of certain metrics.
Can we establish a similar robot control paradigm that both looks back (i.e., generative) and looks ahead (i.e., predictive)?
Contribution. We propose generative predictive control (GPC), a framework that combines a generative robot policy with a predictive dynamical model of the world (i.e., a world model [22]). GPC contains three modules (cf. Fig. 1):
  • Generative policy training. Given expert demonstrations, GPC uses generative modeling to train a policy that approximates the conditional distribution of expert “action chunks” given past observations.
  • Predictive world modeling. A world model simulates future observations conditioned on action chunks and past observations. One would naturally want to learn the world model using the same expert demonstrations as in policy training. However, such a world model can only simulate dynamics of the expert policy and thus is not effective for correcting policy error (see §IV). We show that adding “random exploration” data—cheaper to collect—enriches the data distribution and yield a world model that strengthens the policy at test time. For state-based world modeling, a simple MLP is sufficient to learn accurate dynamics. For vision-based world modeling, we capitalize on conditional video diffusion [25, 6] to predict action-conditioned future observations.
  • Online planning. Inspired by classical model predictive control, and assuming a reward model is available, GPC uses two lightweight online algorithms to strengthen the generative policy with the predictive world model. The first one, GPC-RANK, samples many action proposals from the policy, unrolls the future conditioned on each proposal, and selects the proposal with highest reward. The second one, GPC-OPT, treats the action proposal from the policy as a “warmstart” and optimizes it to maximize reward using gradient ascent, a technique known as single shooting [16]. These two algorithms can be easily combined to perform gradient-based optimization on multiple action proposals (warmstarts) and select the best one.
We evaluate GPC in both simulated and real-world, state-based and vision-based robotic manipulation tasks, demonstrating significant improvement over pure behavior cloning.
Novelty. Among the three modules of GPC, the first one is not new and has become the common practice in robot policy learning [15, 52]. The novelty of GPC lies in learning a predictive world model—crucially, using random exploration data—to strengthen the generative policy with online planning (as opposed to, e.g., offline RL [23, 28, 56, 17, 44]). Moreover, different from previous works leveraging generative models for world modeling, GPC learns a world model that is explicitly action-conditioned. In other words, the world model is “controllable” by the robot’s actions. We show such a controllable world model enables robust real-world robotic manipulation.
Paper organization. We formalize the overall pipeline of GPC in §II and detail the approaches for state-based and vision-based world modeling in §III. We present experimental results in §IV and conclude in §VI. Review of related works and how GPC differs from them are described in §V.
II Overview of Generative Predictive Control
Let $I_{t}$ be the information vector summarizing the state of the robot and the environment up to time step $t$. For state-based robot control, $I_{t}:=x_{t}$ is the (low-dimensional) state of the robot and its environment (e.g., poses); for vision-based robot control, $I_{t}:=o_{t-H:t}$ is the history of (high-dimensional) visual observations where each $o_{t}$ is a single (or multi-view) image(s). Let $a_{t}$ be the robot action at time $t$ and $a_{t:t+T}$ be an action chunk of length $T+1$. Our goal is to design a robot control policy that decides $a_{t:t+T}$ based on $I_{t}$ to solve certain tasks (e.g., described by language instructions).
Let us formalize the three modules of GPC.
Generative policy training. We ask humans to teleoperate robots to solve the tasks. The expert demonstrations are then divided into clips of data containing pairs of state information and action chunks, i.e., a dataset ${\cal D}_{\text{expert}}^{P}=\{I_{t}^{i}\leftrightarrow a_{t:t+T}^{i}\}_{i=1}^{N_{\text{expert}}}$ with number of samples $N_{\text{expert}}$. Policy learning becomes a supervised learning problem with input $I_{t}$ and output $a_{t:t+T}$. [^1] It is also common to include language instructions as input to the policy as in vision-language-action (VLA) models [32]. GPC can generalize to VLAs in a straightforward way. We exclude language here to focus on motion planning. In GPC, we follow the diffusion policy learning framework [15]. Specifically, we feed $I_{t}$ into a network that parametrizes the score function of the conditional distribution $p(a_{t:t+T}|I_{t})$, using which Gaussian noise vectors are gradually diffused to expert action chunks. The score function network is trained using denoising prediction [24]. Policy learning leads to a policy network ${\cal P}(\cdot)$ that predicts action chunks $a_{t:t+T}$ given state information $I_{t}$. It is important to note that ${\cal P}(\cdot)$ is stochastic and effectively a sampler from the (approximate) conditional distribution $p(a_{t:t+T}|I_{t})$, meaning that given $I_{t}$, we can draw many possible action chunks $(a_{t:t+T}^{(1)},\dots,a_{t:t+T}^{(K)})$ from ${\cal P}(\cdot)$. We call each action chunk $a_{t:t+T}$ an action proposal.
Predictive world modeling. The fact that ${\cal P}(\cdot)$ is stochastic leads to a natural question: which one of the $K$ action proposals should be selected? The world model precisely aims to answer this question by predicting the future consequences of each action proposal. To build the world model, we need pairs of $(I_{t},a_{t:t+T})$ (current state information plus an action chunk) and $I_{t+1:t+T+1}$ (future state information). The expert demonstrations used for policy training already contain such pairs, i.e., a dataset ${\cal D}_{\text{expert}}^{W}=\{(I^{i}_{t},a^{i}_{t:t+T})\leftrightarrow I_{t+1:t+T+1}^{i}\}_{i=1}^{N_{\text{expert}}}$. Nevertheless, as we show in §IV, if we train the world model only using ${\cal D}_{\text{expert}}^{W}$, the predictive power of the world model is insufficient to guide action proposal selection. One can collect more expert demonstrations to enlarge the datasets, but it can be expensive. Instead, we posit that random exploration data are not only easier to collect, but also rich in dynamics. Therefore, we collect another exploration dataset ${\cal D}^{W}_{\text{explore}}=\{(I^{i}_{t},a^{i}_{t:t+T})\leftrightarrow I_{t+1:t+T+1}^{i}\}_{i=1}^{N_{\text{explore}}}$ where humans (or any other controllers) are asked to randomly perturb the system without solving any tasks. In control theory terminology, the goal of this is to perform “system identification” with “sufficient excitation” [37]. We combine the two datasets ${\cal D}^{W}:={\cal D}^{W}_{\text{expert}}\cup{\cal D}^{W}_{\text{explore}}$ to train the world model. Details of the design of the world model architecture are deferred to §III. For now, the reader can assume we obtain a model ${\cal W}(\cdot)$ that predicts the future $I_{t+1:t+T+1}$ conditioned on the current state information and the action proposal $(I_{t},a_{t:t+T})$.
Online planning. We now formalize the two online planning algorithms, namely GPC-RANK and GPC-OPT, that combine the policy ${\cal P}(\cdot)$ and the world model ${\cal W}(\cdot)$ (cf. Fig. 1). We assume a reward model ${\cal R}(\cdot):I_{t+1:t+T+1}\rightarrow r_{t}$ is available that predicts the reward given the future state information, e.g., a small neural network that is separately trained.
  1. (i)
    GPC-RANK is based on the simple intuition to pick the action proposal with highest reward. Formally, we sample $K$ action proposals from the policy
    $(a_{t:t+T}^{(1)},\dots,a_{t:t+T}^{(k)},\dots,a_{t:t+T}^{(K)})\sim{\cal P}(I_{t}),$ (1)
    and output the best proposal by passing them through the world model ${\cal W}(\cdot)$ and computing respective rewards:
    $\pi(I_{t})=a_{t:t+T}^{(k_{\star})},\quad k_{\star}\in\operatorname*{arg\,max}_{k=1,\dots,K}\ {\cal R}({\cal W}(I_{t},a_{t:t+T}^{(k)})),$ (2)
    where we used the notation $\pi(\cdot)$ to denote the online policy in contrast to the offline policy ${\cal P}(\cdot)$. The underlying assumption of GPC-RANK is that, due to distribution shifts, the optimal action chunk at test time may not have the highest likelihood in ${\cal P}(\cdot)$, and thus may need more samples to be discovered. The world model—assuming it is learned from a richer distribution because of random exploration—will pick up the optimal action chunk through (2). A nice property of GPC-RANK is that it can be easily parallelized.
  2. (ii)
    GPC-OPT takes a different perspective by directly solving the reward maximization problem given the world model:
    $\max_{a_{t:t+T}}{\cal R}({\cal W}(I_{t},a_{t:t+T})),$ (3)
    where the action chunk are viewed as decision variables. The difficulty is, however, problem (3) can be highly nonconvex due to the nonlinearity and nonsmoothness of ${\cal W}(\cdot)$ and ${\cal R}(\cdot)$. Fortunately, the pretrained policy ${\cal P}(\cdot)$ comes to rescue because it narrows down the search space by providing good initializations to solve (3), i.e., warmstarts. Formally, we sample an action chunk $\hat{a}_{t:t+T}$ from ${\cal P}(\cdot)$ and perform $M$ steps of gradient ascent starting from $\hat{a}_{t:t+T}$ (assuming ${\cal R}(\cdot)$ is differentiable):
    $\hat{a}_{t:t+T}^{(0)}:=\hat{a}_{t:t+T}\sim{\cal P}(I_{t}),\text{ and for }\ell=1,\dots,M\text{ do:}$:
    $\hat{a}_{t:t+T}^{(\ell)}=\hat{a}_{t:t+T}^{(\ell-1)}+\eta^{(\ell)}\nabla_{a_{t:t+T}}{\cal R}({\cal W}(I_{t},\hat{a}_{t:t+T}^{(\ell-1)})),$ (4)
    where $\eta^{(\ell)}>0$ are step sizes. The online policy outputs the final optimized action chunk
    $\pi(I_{t})=\hat{a}_{t:t+T}^{(M)}.$ (5)
    Since the gradients in ((ii)) are readily available from automatic differentiation, gradient ascent can be carried out using off-the-shelf optimizers like ADAM. Readers familiar with the optimal control literature will recognize GPC-OPT is the classical single shooting algorithm [16]. Compared to GPC-RANK which evaluates the world model $K$ times, GPC-OPT evaluates the world model $M$ times. The former can be parallelized but the latter cannot.
GPC-RANK and GPC-OPT can be combined. We can sample $K$ action proposals from ${\cal P}(\cdot)$ and for each proposal perform reward maximization according to ((ii)), leading to $K$ optimized action chunks. We pick the optimized action chunk with highest reward. GPC-RANK+OPT can be viewed as solving the optimization (3) from multiple initializations [50], and requires $K\times M$ evaluations of the world model.
We provide a succinct summary of GPC in Algorithm 1.
Input: Expert action demonstrations ${\cal D}^{P}_{\text{expert}}=\{I_{t}^{i}\leftrightarrow a^{i}_{t:t+T}\}_{i=1}^{N_{\text{expert}}}$; Expert trajectories and random exploration trajectories ${\cal D}^{W}_{\text{expert}}\cup{\cal D}^{W}_{\text{explore}}=\{(I_{t}^{i},a^{i}_{t:t+T})\leftrightarrow I^{i}_{t+1:t+T+1}\}_{i=1}^{N_{\text{expert}}+N_{\text{explore}}}$; Reward model ${\cal R}$; Positive integers $K$, $M$
1
2// Generative policy training
[…]
4
5// Predictive world modelingIII)
[…]
7
8// Online planning given $I_{t}$
9for $k=1,\dots,K$ do
10       // Sample action proposal
11      $a^{(k)}_{t:t+T}\sim{\cal P}(I_{t})$;
12
13      // Maximize reward
14      Set $\hat{a}^{(0)}_{t:t+T}=a^{(k)}_{t:t+T}$;
15
16      for $\ell=1,\dots,M$ do
17             $\hat{a}^{(\ell)}_{t:t+T}=\hat{a}^{(\ell-1)}_{t:t+T}+\eta^{(\ell)}\nabla_{a_{t:t+T}}{\cal R}({\cal W}(I_{t},\hat{a}_{t:t+T}^{(\ell-1)}))$;
18      Set $a^{(k)}_{t:t+T}=\hat{a}^{(M)}_{t:t+T}$;
19
20
21// Rank optimized action proposals
22Find $k_{\star}\in\operatorname*{arg\,max}_{k=1,\dots,K}\ {\cal R}({\cal W}(I_{t},a_{t:t+T}^{(k)}))$ ;
23
24Return $a_{t:t+T}^{(k_{\star})}$ ;
25
Algorithm 1 Generative Predictive Control (GPC)
Remark 1 (Connection with Offline RL).
It is worth noting that the composition of the reward model ${\cal R}(\cdot)$ and the world model ${\cal W}(\cdot)$ can be identified as the so-called Q-value function in offline RL, as they predict the reward of an action chunk $a_{t:t+T}$ given the current information $I_{t}$. However, there are three subtle differences between GPC and offline RL. (a) GPC learns an explicit dynamics model instead of just approximating the Q-value function (this is similar to model-based offline RL). (b) The composition of ${\cal R}(\cdot)$ and ${\cal W}(\cdot)$ provides a Q-value function for “action chunks” while the usual Q-value function involves a single action. (c) GPC learns the dynamics model using diverse trajectories (without rewards), instead of only expert demonstrations. In §IV, we compare GPC with an offline RL method adapted from single action to action chunk and show that GPC outperforms offline RL.
The above overview of GPC highlights the importance of the world model in both ranking and optimizing action proposals. Indeed, the concept of a learned world model dates back to early pioneering works on deep visual foresight [20]. In the next section, we will show that modern diffusion models are the key enabler for learning (near) physics-accurate visual world models that were not possible before.
Figure 2: Illustration of planar pushing.
Figure 3: Diffusion-based Visual World Modeling. Our visual world model consists of a sequence of recursive single-step image predictors. Each single-step image predictor is designed as a conditional diffusion process.
III World Model Learning
Recall that the world model takes as input $(I_{t},a_{t:t+T})$ and outputs $I_{t+1:t+T+1}$. We will briefly present the world model design when $I_{t}=x_{t}$ contains low-dimensional states (§III-A) and focus on a diffusion-based world model when $I_{t}=o_{t-H:t}$ contains high-dimensional images (§III-B).
Planar pushing. Before getting started, since we will focus on planar pushing in our experiments §IV, we briefly introduce the problem setup. As shown in Fig. 2, the planar pushing task involves pushing an object (e.g., a T block) to align with the target using a pusher. We assume the target position is fixed and known, and hence we do not need to predict it. The state of the object is fully described by its 2D coordinates $p=(p_{x},p_{y})$ and 1D angle $\theta$. The action space is $a\in\mathbb{R}^{2}$ containing pushing velocities in the $x$ and $y$ directions. We choose the planar pushing task for two reasons. First, it is easy to understand and simple to visualize. Second, it is also a long-standing difficult task because it is contact-rich. Depending on whether the pusher makes or breaks contact and whether the contact is sticking or sliding, the dynamics of the task involves multiple different modes and requires intricate physics reasoning [26]. This makes planar pushing a both elementary and challenging task for world model learning.
III-A State-based World Modeling
For state-based world modeling, with $x=(p,\theta)$ being the state of the object, we learn a single-step dynamics model:
$x_{t+1}=f_{\text{state}}(x_{t},a_{t}).$ (6)
As pointed out by [64], when rotations are involved in machine learning, it is crucial to choose a rotation representation that is continuous. For example, the angle representation $\theta$ is discontinuous due to the “wrap $2\pi$” issue. Therefore, we augment the state as
$x=(p,\cos(\theta),\sin(\theta))$ (7)
for both the input and output of $f_{\text{state}}$ in (6) (we normalize the output of $f_{\text{state}}$ to satisfy $\sin^{2}(\cdot)+\cos^{2}(\cdot)=1$). We also found it beneficial to have two separate networks, one predicting the position and the other predicting the rotation. In our implementation, for each of the network, we used a six-layer MLP with hidden dimension $128$.
During testing, we need to predict $T$ steps into the future given a sequence of actions $a_{t:t+T}$. We do so by recursively applying the single-step predictor (6). As we will show in §IV, this simple MLP-based single-step predictor is sufficient for robust online planning.
III-B Vision-based World Modeling
State-based world modeling requires accurate state estimation, which is possible in a lab environment with infrastructure such as AprilTags [46] or motion capture but can be challenging in the open environment. Therefore, it is desired to directly learn visual dynamics, i.e., a visual world model.
Diffusion-based visual world modeling. Motivated by the success of using diffusion models for image generation, we design a visual world model based on conditional diffusion. Recall that the input to the world model is $I_{t}=o_{t-H:t}$ (the sequence of past images) and $a_{t:t+T}$, and the output is $I_{t+1:t+T+1}:=o_{t+1:t+T+1}$ (a sequence of future images). Similar as the state-based world modeling, we design the visual world model as a recursive application of single-step image predictors. We denote the single-step predictor as
$o_{t+1}=f_{\text{vision}}(o_{t-H:t},a_{t}),$ (8)
where $f_{\text{vision}}$ is a conditional diffusion model. The unique property of diffusion is that $o_{t+1}$ is not generated in a single feedforward step, but rather through a sequence of denoising steps. In particular, let $o_{t+1}^{{\cal T}}$ be drawn from a white Gaussian noise distribution, then $f_{\text{vision}}$ proceeds as:
$o_{t+1}^{\tau-1}=D_{\phi}(o_{t+1}^{\tau},\tau,o_{t-H:t},a_{t}),\tau={\cal T},\dots,1,$ (9)
where ${\cal T}$ is the total number of denoising steps and the output of $f_{\text{vision}}$ is $o_{t+1}=o_{t+1}^{0}$. In (9), $\tau$ is the denoising step index and $D_{\phi}$—the “denoiser”—is a neural network with weights $\phi$. We use the same architecture for $D_{\phi}$ as in [5], which contains convolutions, action embedding, and a U-Net, as illustrated in Fig. 3. The denoiser $D_{\phi}$ is trained by adding random noises to the clean images and then predicting the noise.
Two-phase training. To improve the accuracy and consistency of visual world modeling, we train it in two phases. In phase one, we train only the single-step image predictor (8), i.e., with supervision only on a single image $o_{t+1}$. In phase two, we recursively apply the single-step predictor (8) for $T$ times to obtain a sequence of future images $o_{t+1:t+T+1}$, and jointly supervise them with groundtruth images.
We use observation horizon $H=4$ in the visual world modeling, and ${\cal T}=3$ diffusion steps in $f_{\text{vision}}$.
Remark 2 (Freeze the Noise).
It should be emphasized that the single-step image predictor $f_{\text{vision}}$ in (8) is “stochastic” where the stochasticity comes from the randomness in the initial noise $o_{t+1}^{{\cal T}}$. In classical generative modeling, stochasticity is desired because we want the model to generate diverse images. However in our paper, we want to use the diffusion-based world model for control. That is, we want to understand the impact of the “action” on the future, rather than the impact of the random “noise” on the future. Therefore, we freeze the noise $o_{t+1}^{{\cal T}}$ to be zero at inference time, which makes the world model deterministic—outputing the most likely future predictions (because $o_{t+1}^{{\cal T}}=0$ attains the peak probability in the Gaussian distribution). In fact, without freezing the noise, GPC-OPT does not work because the gradient of the reward with respect to the actions (cf. ((ii))) would be “contaminated” by the noise and thus does not guide online planning.
IV Experiments
In this section, we focus on evaluation of the GPC framework on planar pushing and comparison to baseline behavior cloning. As mentioned in §III, we choose planar pushing because it is an elementary but also quite challenging task due to the complicated contact and collision dynamics involved.
IV-A State-based Planar Pushing in Simulation
We first study the setup where the groundtruth state of the object (i.e., $(p_{x},p_{y},\theta)$ is available through a simulator. We use the same simulator as in [15] with a T-shape object.
Policy. We first train a state-based diffusion policy ${\cal P}(\cdot)$ [15] from which we can draw action proposals, as explained in §II. We use $500$ expert demonstrations to train the policy network for $300$ epochs with AdamW optimizer of learning rate $10^{-4}$ and weight decay $10^{-6}$. We use observation horizon $H=2$, prediction horizon $T=16$, and action horizon $8$. [^2] This means the policy is executed in a receding horizon fashion. The policy predicts $16$ steps of future actions and only $8$ of them are executed in the environment, which creates implicit feedback.
World model. Then we need the world model to predict the future consequences of each action proposal. As in §III-A, we use MLP networks to build a single-step state predictor and recursively apply it $T$ times to form the world model ${\cal W}(\cdot)$. As introduced in §II, to ensure the dynamics of the world model is learned from a richer distribution than the expert demonstrations, the training dataset of the world model is expanded with random exploration demonstrations which attempt to record more pusher-object interactions at different contact points. To be specific, we let human operators (who are different from the operator of the expert demonstrations) collect $7$ long random play demonstrations and 1 special long demonstration that randomly collides the object with the boundary wall. In addition, we also generate $7$ sets of “noisy expert demonstrations” where the original expert actions are contaminated with different scales of Gaussian noise.
Reward. After obtaining the predicted consequences of the action proposals (i.e., predicted future states of $p_{x},p_{y},\theta$ of the object), we define the reward ${\cal R}(\cdot)$ using a registration loss between the target and the object. Particularly, the registration loss computes the sum of the squared distances between the vertices of the object and the target. Clearly, this reward function is differentiable with respect to the predicted states.
Online planning. Now we can use the policy ${\cal P}(\cdot)$, the world model ${\cal W}(\cdot)$, and the reward model ${\cal R}(\cdot)$ for online planning, as prescribed in the GPC Algorithm 1. Specifically, In GPC-RANK, we pick the action chunk with highest reward (smallest registration loss) from the $K$ action proposals. In GPC-OPT, we sample one action chunk from ${\cal P}(\cdot)$ and perform $M$ steps of gradient ascent to optimize this action chunk for a higher reward (smaller registration loss). In GPC-RANK+OPT, we sample $K$ action proposals, perform $M$ gradient ascent steps for each of the action proposal, and finally pick the optimized action chunk with the highest reward.
Questions. We hope to answer the following questions through our experiments:
  • How much does GPC outperform the state-of-art behavior cloning method (e.g., diffusion policy) and how do $K$ and $M$ affect the performance?
  • How does GPC compare with Offline RL based method?
  • What is the effect of training the world model using both expert and random exploration data?
Results. In Table I, we study how different variations of $K$ and $M$ affect the performance. When $K=1$ and $M=0$, this is the baseline state-based diffusion policy trained from behavior cloning [15]. In GPC-RANK, if we increase the number of action proposals to $K=100$, the score increases to around 0.9, which is about $10\%$ increase from the baseline. Notably, due to parallelization inside GPC-RANK, the runtime does not increase when $K=100$. We also tried larger $K$, such as $1000$ and $5000$, and the performance further increases to $0.934$. Note that in the last column of Table I, we provided a “performance upper bound” which is obtained by using the groundtruth simulator (instead of the learned world model) to rank action proposals at $K=5000$. For GPC-OPT, if we sample one action proposal and perform gradient ascent from it (e.g., doing $30$ gradient steps), the score can be increased to $0.886$, which is also quite significant compared to the baseline. Then in GPC-RANK+OPT, we tried different combinations of $K$ and $M$. The improvement remains substantial, indicating the robustness of GPC and shows both sampling multiple proposals and gradient-based optimization consistently improve the performance.
In the first two columns of Table I, we present the results for another baseline of using Offline RL to rank and select actions. The Offline RL method follows that of [44] and is described in Appendix -A. Crucially, the Offline RL method only uses the expert demonstrations and does not leverage the random exploration data (because the random exploration data do not have labeled rewards). We can see that the performance of Offline RL is similar or even worse than the case of behavior cloning. A plausible reason for this is that the Offline RL method does not explicitly seek to learn the dynamics from the random explorations and hence the learned value function may be misleading for ranking action proposals.
Finally, in Table II, we present the results of using a world model trained without random exploration data. We observe that, when the world model is only trained from expert demonstrations, the performance of both GPC-RANK and GPC-OPT gets worse. This demonstrates the importance of using random exploration data to learn an accurate world model.
Offline RL GPC-RANK GPC-OPT GPC-RANK+OPT With GT simulator $K=100$ $K=5000$ $\begin{subarray}{c}K=1\\ M=0\end{subarray}$ $\begin{subarray}{c}K=100\\ M=0\end{subarray}$ $\begin{subarray}{c}K=1000\\ M=0\end{subarray}$ $\begin{subarray}{c}K=5000\\ M=0\end{subarray}$ $\begin{subarray}{c}K=1\\ M=30\end{subarray}$ $\begin{subarray}{c}K=10\\ M=30\end{subarray}$ $\begin{subarray}{c}K=20\\ M=30\end{subarray}$ $\begin{subarray}{c}K=30\\ M=20\end{subarray}$ Score $0.788$ $0.745$ $0.812$ $0.898$ $0.932$ $0.934$ $0.886$ $0.912$ $0.914$ $0.891$ $0.952$
TABLE I: State-based Planar Pushing (Push-T) in Simulation. Score is measured by the IoU metric averaged over $100$ evaluation seeds. $K$ is the number of action proposals sampled from ${\cal P}(\cdot)$ while $M$ is the number of gradient steps performed to maximize reward.
GPC-RANK GPC-OPT $K=100,M=0$ $K=1000,M=0$ $K=5000,M=0$ $K=1,M=30$ Score (using world model trained without random exploration) $0.832$ $0.838$ $0.816$ $0.752$ Score (using world model trained with random exploration) $0.898$ $0.932$ $0.934$ $0.886$
TABLE II: The importance of random exploration in world model learning (state-based Push-T).
IV-B Vision-based Planar Pushing in Simulation
Figure 4: Comparison of different visual world modeling techniques on two examples.
We then proceed to vision-based planar pushing with the same T-shape object.
Policy. We train a vision-based diffusion policy ${\cal P}(\cdot)$ [15] using ResNet18 as the vision encoder and UNet as the action diffusion network. We used $500$ expert demonstrations to train the policy for $300$ epochs with AdamW optimizer of learning rate $10^{-4}$ and weight decay $10^{-6}$. We use observation horizon of 2, prediction horizon of 16, and action horizon of 9.
World model. As in §III-B, we use a diffusion model to build a single-step image predictor and recursively apply it $T$ times to form the world model ${\cal W}(\cdot)$. Same as in §IV-A, the training data includes random explorations.
Reward. Different from the state-based planar pushing, since the future predictions are images, we cannot directly compute a reward signal. Therefore, we first train a pose prediction network with ResNet18 plus MLP, and then compute the registration loss using the predicted object pose. Note that this learned reward model ${\cal R}(\cdot)$ is still differentiable.
Online planning: With the vision-based policy, the visual world model, and the learned reward model, we follow the same online planning protocol described in §IV-A.
Quality of visual world modeling. Before presenting the results of GPC for online planning, we first study whether our visual world modeling technique is better than some baselines. Particularly, we chose two baselines, one from the pioneering deep visual foresight [20], and the other from the more recent AVDC paper [19]. The world model in deep visual foresight leverages CNN and LSTM for future prediction, and we follow the Github implementation. [^3] https://github.com/Xiaohui9607/physical_interaction_video_prediction_pytorch AVDC is originally designed for video diffusion conditioned on language, and we modify its official implementation [^4] https://github.com/flow-diffusion/AVDC to condition on robot actions. Notably, AVDC is a multiple-step world model as it simultaneously diffuses future images together, unlike our recursive structure. Fig. 4 compares the quality of future predictions of three methods (ours, deep visual foresight, and AVDC) against the groundtruth in two examples. We see that diffusion-based visual world modeling (ours and AVDC) clearly dominates the classical deep visual foresight. Between ours and AVDC, we can observe that the future predictions from our method are better aligned with the groundtruth.
Planning performance. Same as §IV-A, we study the impact of $K$ and $M$ on the performance of GPC, its comparison with Offline RL, and the importance of random exploration.
In Table III, we study how different variations of $K$ and $M$ affect the performance of GPC. We can see that (a) GPC-RANK increases around $15\%$ from the diffusion policy baseline (when $K=1,M=0$); (b) GPC-OPT increases around $25\%$ from the diffusion policy baseline; and (c) GPC-RANK+OPT increases up to $40\%$ from the baseline. Contrasting the vision-based results with the state-based results, we observe that GPC-OPT appears to have a larger advantage in vision-based push-T. This indeed makes sense because the baseline vision-based diffusion policy only has a score of $0.642$, leaving a lot of room for optimization to find better actions.
In the first column of Table III, we show the results of Offline RL. Unfortunately, Offline RL still does not perform as well as even the baseline vision-based diffusion policy.
In Table IV, we compare the performance of GPC-RANK and GPC-OPT when trained with or without random exploration. Consistent with our observations in IV-A, without training on random exploration, the performance of online planning significantly decreases.
Offline RL GPC-RANK GPC-OPT GPC-RANK+OPT $K=100$ $K=1,M=0$ $K=50,M=0$ $K=100,M=0$ $K=200,M=0$ $K=1,M=25$ $K=5,M=10$ $K=10,M=25$ Score $0.620$ $0.642$ $0.698$ $0.739$ $0.719$ $0.791$ 0.824 0.882
TABLE III: Vision-based Planar Pushing (Push-T) in Simulation. Score is measured by the IoU metric averaged over $100$ evaluation seeds. $K$ is the number of action proposals and $M$ is the number of gradient steps.
GPC-RANK GPC-OPT $K=100,M=0$ $K=1,M=25$ Score (using world model trained without random exploration) $0.647$ 0.652 Score (using world model trained with random exploration) $0.739$ $0.791$
TABLE IV: The importance of random exploration in world model learning (vision-based push-T).
IV-C Real-world Vision-based Planar Pushing
We then test GPC in real-world planar pushing. We not only show GPC is effective in “plain push-T”, but also show it can work for push-T with collision objects.
Figure 5: Real-world setup.
Setup. Fig. 5 depicts our real-world setup for vision-based planar pushing. The pusher is attached to the end-effector of a Franka Panda robotic arm to push a T block into the target position. An overhead Intel RealSense camera is used to provide image observations to the pusher. To train the reward model, we need to obtain block information as training data so we attached three AprilTags [46] to the T-block, which enables accurate pose estimation from image frames even in the case of occlusion. To collect real-world expert demonstrations, we follow Chi et al. [15] and use a SpaceMouse to teleoperate the robotic arm to push the T block. A and R blocks are the objects that T may collide into during pushing.
IV-C1 Plain push-T
We first test on plain push-T (i.e., without collision with other objects). We train the policy with $100$ expert push-T demonstrations using the architecture mentioned in §IV-B. Because of the slower moving speed of real-world pusher, we use a larger prediction horizon $T=32$. Same as §IV-B, we train the world model using all expert demonstrations and a few random explorations. For the reward model, we rely on AprilTag [46] to obtain the ground-truth object poses and use them to train the reward model ${\cal R}(\cdot)$. To be clear, our GPC pipeline in real world is still purely vision-based and the AprilTags are only used to obtain training data for the reward model (training the pose predictor). We follow the same online planning protocol described in §IV-A. We test GPC-RANK and GPC-OPT on a set of 10 push-T tasks where the T block is initialized at positions not seen during training.
Performance. While the baseline model succeeds 5 out of 10 times, GPC-RANK succeeds 7 out of 10 times and GPC-OPT succeeds 7 out of 10 times. Fig. 10(a) shows the trajectories for one test. More testing trajectories are shown in Appendix -B and supplementary videos.
Refer to captionRefer to caption
(a) World model prediction in GPC-RANK
Refer to captionRefer to caption
(b) World model prediction in GPC-OPT
Figure 6: World model prediction in GPC for push-T. [Top] For GPC-RANK, we show the action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$. Every row corresponds to the set of image consequences for an action proposal. We only pick $6$ action proposals and downsample the image prediction sequences for displaying purpose. [Bottom] For GPC-OPT, we show optimized action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$. Every row corresponds to the set of image consequences for an optimized action proposal at optimization step $i$. We only pick $6$ optimization steps and downsample the image prediction sequences for displaying purpose.
Quality of world model predictions. Fig. 6 shows the quality of world model prediction of push T for GPC-RANK and GPC-OPT. The images shown are all predicted by the world model. They not only have very good visual effect and look realistic, but also they can accurately predict the dynamics reaction from given actions. From the last column, we can see even when the differences between action sequences are very small, the world model can still capture the details and predict the reaction in a meticulous way. This is very important to make GPC work.
IV-C2 Push-T under collision
Next, we hope to investigate how GPC works under more complex scenarios, involving harder tasks and more complex dynamics. We put an object of shape “A” on top of the target position so that when pushing T, it will inevitably collide with object A. The training data is $100$ push-T expert demos with collision of A and some random exploration demos.
Performance. We test $5$ seeds in total, and baseline model, GPC-RANK, GPC-OPT succeed for $2$, $4$, $3$ seeds respectively.
We then proceed to test in an even harder situation: we put objects of shape “A” and “R” on top of the target position at the same time. Case 1: When pushing T, T may collide with A and R at the same time. Case 2: When pushing T, T would collide with one letter (e.g., “R”) and this letter will collide with another letter (e.g., “A”). We train the policy and world model here using $100$ plain push-T expert demonstrations, $100$ push-T colliding A expert demonstrations, and $100$ push-T colliding A&R expert demonstrations.
Performance. In the total of $5$ testing seeds, baseline model, GPC-RANK, GPC-OPT succeed $2$, $3$, $4$ respectively.
Fig. 10(b)(c) show the trajectories for both of these tests. More testing trajectories are shown in Appendix -B and supplementary videos.
Refer to captionRefer to caption
(a) World model prediction in GPC-RANK
Refer to captionRefer to caption
(b) World model prediction in GPC-OPT
Figure 7: World model prediction in GPC for push-T with collision of A. [Top] For GPC-RANK, we show the action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$. Every row corresponds to the set of image consequences for an action proposal. We only pick $6$ action proposals and downsample the image prediction sequences for displaying purpose. [Bottom] For GPC-OPT, we show optimized action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$. Every row corresponds to the set of image consequences for an optimized action proposal at optimization step $i$. We only pick $6$ optimization steps and downsample the image prediction sequences for displaying purpose.
Refer to captionRefer to caption
(a) World model prediction in GPC-RANK
Refer to captionRefer to caption
(b) World model prediction in GPC-OPT
Figure 8: World model prediction in GPC for push-T with collision of A and R. [Top] For GPC-RANK, we show the action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$. Every row corresponds to the set of image consequences for an action proposal. We only pick $6$ action proposals and downsample the image prediction sequences for displaying purpose. [Bottom] For GPC-OPT, we show optimized action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$. Every row corresponds to the set of image consequences for an optimized action proposal at optimization step $i$. We only pick $6$ optimization steps and downsample the image prediction sequences for displaying purpose.
Quality of world model predictions. Fig. 7 and Fig. 8 show the quality of world model prediction for GPC-RANK and GPC-OPT, having collisions during push-T. The images shown are all predicted by the world model. With adding some collision objects, the world model is not only able to give near perfect prediction (i.e. no deformation) for a complex scene composing several objects, but also it can accurately predict the dynamics among these objects.
IV-C3 Generalization to unseen cases
Last but not least, we want to understand whether GPC can help when generalizing to unseen cases. We use the policy and world model trained with push-T, push-T colliding A, push-T colliding A&R expert demos. But we perform the test to push T when an object of letter “R” is on top of the target. This situation has never been seen in any of the training data. We test three different starting positions for T and R.
Performance. GPC-RANK and GPC-OPT succeed in all of the three tests while baseline model fails in one of the cases. The reason behind this is that when generalizing to unseen cases, the policy’s performance degrades. So we need to either generate more samples from the action policy’s distribution (like giving more opportunities to generate an optimal proposal) or do gradient ascent to optimize the action proposal.
Fig. 11 shows one test. More testing trajectories are shown in Appendix -B and supplementary videos.
Refer to captionRefer to caption
(a) World model prediction in GPC-RANK
Refer to captionRefer to caption
(b) World model prediction in GPC-OPT
Figure 9: World model prediction in GPC for push-T with collision of R. [Top] For GPC-RANK, we show the action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$. Every row corresponds to the set of image consequences for an action proposal. We only pick $6$ action proposals and downsample the image prediction sequences for displaying purpose. [Bottom] For GPC-OPT, we show optimized action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$. Every row corresponds to the set of image consequences for an optimized action proposal at optimization step $i$. We only pick $6$ optimization steps and downsample the image prediction sequences for displaying purpose.
Quality of world model predictions. Fig. 9 shows the quality of world model prediction of push T under collision in unseen scenarios for GPC-RANK and GPC-OPT. A scenario only composes of T and R is never seen in the training data. However, our world model can still give good predictions for action sequences here. It verifies the ability of generalization for our world model.
Remark 3 (Hallucination).
One of the important issues of generative modeling is hallucination, and our visual world model is no exception. In particular, there are several cases where the world model predicts images of the future that violates rigid-body physics. For example, the objects T, A, and R may penetrate each other, as we show in Appendix -C. Interestingly, even though the predictions can violate physics, using such predictions for planning still leads to GPC’s superior performance. We believe learning a visual world model that strictly satisfies the laws of physics remains a challenging task, and strategies such as data scaling and baking in physics priors may help solve this challenge.
(a) Test for push-T
(b) Test for push-T with collision of A
(c) Test for push-T with collision of A and R
Figure 10: Real-world tests for plain push-T and push-T with collision. In every test, top row shows trajectories of baseline model ($K=1,M=0$), middle row shows trajectories of GPC-RANK ($K=10,M=0$), and last row shows trajectories of GPC-OPT ($K=0,M=25$). The first and last column show the initial and final position of the object, respectively. The middle five columns show overlaid trajectories generated by the policies. Results for other real-world tests are shown in Appendix -B and supplementary videos.
Figure 11: Real-world test for out-of-domain scene. In every test, top row shows trajectories of baseline model ($K=1,M=0$), middle row shows trajectories of GPC-RANK ($K=10,M=0$), and last row shows trajectories of GPC-OPT ($K=0,M=25$). This displays GPC would help to make policy succeed in unseen scene while baseline failed. Remaining results for push-T with collision of R are shown in Appendix -B and supplementary videos.
V Related Work
Generative modeling for robotics. A large body of recent work has explored how generative models can be integrated with robotics [52]. Generative models have been used extensively to represent policies [43, 42, 27, 15, 11, 7, 36, 29], where models are often trained using large datasets of task demonstrations. In addition, generative models have also been used as a source of generating additional data [55, 54, 14] as well as models of the dynamics of the world [40, 12, 60].
Finally, a large body of work has explored how generative models can be used for planning [28, 3, 59, 13, 41]. Such methods directly train a generative model over sequences of state and action trajectories, and use the generation process to optimize actions. Different from these works, our work proposes a generative and predictive control (GPC) framework that combines a generative policy with a predictive world model through online planning to obtain more robust policy execution on new instances of a task. The modular design of GPC also allows for training foundation policies and foundation world models separately, using different data.
Visual world modeling. Learning predictive visual models of the world has a long history [58, 45, 22, 47]. Recently, video generative models have emerged as a powerful tool for modeling the physical world [19, 60, 39, 33, 53, 9, 8]. Such models have been used to initialize policies [19, 33, 9], as interactive simulators [60, 39, 8], and integrated with downstream planning [18, 53]. Different from these works, our work uses a video model as an action-conditioned dynamics model that enables us to build a generative control framework that online computes a sequence of actions to optimizes a task-specific reward function.
VI Conclusion
We presented generative predictive control (GPC), a learning control paradigm that combines a generative policy with a predictive world model during online planning. It is worth noting that GPC bears great similarity with the classical literature on predictive control in the sense that the dynamics model is first identified and then optimization is performed online to decide actions. The new insights offered by GPC, perhaps, are that (a) modern conditional diffusion models provide a powerful tool to learn controllable action-conditioned visual world models; and (b) a generative policy trained from behavior cloning provides good warmstarts for online optimization that seem promising to partially tackle the nonconvexity of online optimization. Through an in-depth investigation on planar pushing—both simulated and real—we demonstrated the superior robustness of GPC in both state-based and vision-based robotic manipulation.
Limitations. Two issues are worth highlighting. The first issue is already mentioned in Remark 3 and concerns the problem of hallucinating future predictions that violate the laws of physics. To address this, it may be possible to leverage the so-called world foundation models [2]—trained on Internet-scale videos—and finetune them for specific robotic tasks. The second issue is associated with the computational complexity of using an expensive diffusion-based world model in online planning. It is worth noting that, although GPC dominates the baseline in terms of success rates, it is generally slower than the baseline due to predicting the future using the world model. A promising solution is to leverage consistency models [51] and one-step diffusion [61] for faster world model prediction.
Acknowledgments
This work was partially supported by the Harvard University Dean’s Competitive Fund for Promising Scholarship. The authors thank William Zhang for his help on hardware design and prototyping. We thank members of the Harvard Computational Robotics Group for their contributions in collecting demonstrations. Their collective support has been instrumental in the completion of this work.

### v2

@@
 We present generative predictive control (GPC), a learning control framework that (i) clones a generative diffusion-based policy from expert demonstrations, (ii) trains a predictive action-conditioned world model from both expert demonstrations and random explorations, and (iii) synthesizes an online planner that ranks and optimizes the action proposals from (i) by looking ahead into the future using the world model from (ii).
-Crucially, we show that conditional video diffusion allows learning (near) physics-accurate visual world models and enable robust visual foresight.
-Focusing on planar pushing with rich contact and collision, we show GPC dominates behavior cloning across state-based and vision-based, simulated and real-world experiments.
-Figure 1: Generative Predictive Control (GPC).
-Given visual (or state) observations, a generative policy—trained from behavior cloning—proposes an action chunk.
-Conditioned on the action chunk and past observations, a predictive world model—trained from diverse trajectories—predicts the future observations in “imagination”.
-In (a) GPC-RANK, the policy outputs multiple action proposals and for each proposal the world model predicts its future and associated reward.
-The proposal with highest reward in executed.
-In (b) GPC-OPT, the policy outputs a single action proposal and multiple gradient ascent steps are performed to optimize the action proposal for higher reward.
-Both GPC-RANK and GPC-OPT dominate pure behavior cloning.
-Equal contribution
-I Introduction
+Across a variety of robotic manipulation tasks, we demonstrate that GPC consistently outperforms behavior cloning in both state-based and vision-based settings, in simulation and in the real world.
+Figure 1: Generative Predictive Control.
+1 Introduction
 State-of-the-art robot control policies are synthesized using two major paradigms:
-  • Predictive control.
-With access to a predictive dynamical model of the robot and its physical environment, optimal control (OC) and (deep) reinforcement learning (RL) are the workhorses for designing performant controllers in locomotion [57], manipulation [4], and aviation [1, 31].
-While OC typically designs controllers based on analytic dynamical models and online optimization [30], RL can learn a policy via interactions with a complicated and black-box dynamical model [49], either through simulation or in the real world.
-The nature of predictive control is to look ahead, by anticipating how the environment will change and how much reward will be collected.
-  • Generative control.
-Motivated by the success of generative modeling in vision and language, generative control—learning a controller from expert demonstrations using behavior cloning with generative models—is rapidly rising to prominence [15, 52].
-Due to the abundance of demonstrations and the unprecedented distribution learning capacity of generative models, generative control has led to robot foundation policies capable of solving diverse tasks from vision and language [21].
-The nature of generative control is to look back, making decisions based on expert policies in the past.
+Given access to a predictive model of the robot and its environment, predictive control, e.g., optimal control (OC) and reinforcement learning (RL), serves as the primary tool for designing high-performance controllers in locomotion [1], manipulation [2], and aviation [3, 4].
+OC designs controllers using analytic models and online optimization [5], while RL learns policies through interaction with complex, often black-box dynamics [6], either in simulation or the real world.
+At their core, predictive control methods look ahead, anticipating environmental changes and future rewards.
+Inspired by the success of generative modeling in vision and language, generative control—learning controllers from expert demonstrations using behavior cloning with generative models—is rapidly gaining prominence [7, 8].
+Leveraging abundant demonstrations and the powerful distribution learning capabilities of generative models, generative control has enabled the creation of robot foundation policies that can solve a wide range of tasks from vision and language inputs [9, 10].
+At its core, generative control looks back, basing decisions on past expert behaviors.
 Each paradigm has its advantages and limitations.
-When dynamics modeling is accurate and tractable, predictive control delivers precision and robustness (e.g., Boston Dynamics parkour and SpaceX rocket landing).
+When dynamics and reward modeling is accurate and tractable, predictive control delivers precision and robustness (e.g., Boston Dynamics parkour and SpaceX rocket landing).
 However, in many cases (e.g., manipulation with contact [11] and deformable objects [12], vision-based control [13]), the dynamics can be inaccurate or computationally expensive for optimization (as in OC) and interaction (as in RL).
-Generative control, on the other hand, circumvents the need for careful dynamics modeling and directly learns the policy via generative modeling of the expert demonstrations.
+Generative control, on the other hand, circumvents the need for careful dynamics and reward modeling and directly learns the policy via generative modeling of the expert demonstrations.
 However, generative control suffers from distribution shifts (due to its supervised learning nature [14, 9]) and lack of precision and robustness (i.e., they can do lots of tasks but do few of them as well as specialized controllers [15, 16]).
-Moreover, whether the “scaling law” can address these issues is questionable [7, 63, 62].
+Moreover, whether the “scaling law” can address these issues remains unclear [17, 18, 19].
 Humans, nevertheless, look back and look ahead when making decisions.
 We look back at similar past experiences to shortlist a set of proper actions, and look ahead—unroll the environment dynamics in “imagination”—to predict which action plan is optimal in terms of certain metrics.
-Can we establish a similar robot control paradigm that both looks back (i.e., generative) and looks ahead (i.e., predictive)?
-Contribution.
-We propose generative predictive control (GPC), a framework that combines a generative robot policy with a predictive dynamical model of the world (i.e., a world model [22]).
-GPC contains three modules (cf.
-Fig.
-1):
+Can we establish a similar robot control paradigm that both looks back and looks ahead?
+Contribution. We propose generative predictive control (GPC), a framework that combines a generative policy with a predictive world model [20]. GPC consists of three modules (cf. Fig. 1):
   • Generative policy training.
-Given expert demonstrations, GPC uses generative modeling to train a policy that approximates the conditional distribution of expert “action chunks” given past observations.
+Given expert demonstrations, GPC uses generative modeling, particularly diffusion models, to train a policy that approximates the conditional distribution of expert “action chunks” given past observations.
   • Predictive world modeling.
@@
 One would naturally want to learn the world model using the same expert demonstrations as in policy training.
-However, such a world model can only simulate dynamics of the expert policy and thus is not effective for correcting policy error (see §IV).
-We show that adding “random exploration” data—cheaper to collect—enriches the data distribution and yield a world model that strengthens the policy at test time.
+However, such a world model can only simulate dynamics of the expert policy and thus is not effective for correcting policy error (see §4).
+We show that adding “random exploration” data—cheaper to collect—enriches the data distribution and yields a world model that strengthens the policy at test time.
 For state-based world modeling, a simple MLP is sufficient to learn accurate dynamics.
@@
 The first one, GPC-RANK, samples many action proposals from the policy, unrolls the future conditioned on each proposal, and selects the proposal with highest reward.
-The second one, GPC-OPT, treats the action proposal from the policy as a “warmstart” and optimizes it to maximize reward using gradient ascent, a technique known as single shooting [16].
-These two algorithms can be easily combined to perform gradient-based optimization on multiple action proposals (warmstarts) and select the best one.
+The second one, GPC-OPT, treats the action proposal from the policy as a “warmstart” and optimizes it to maximize reward using gradient ascent, in the spirit of single shooting [23].
+These two algorithms can be combined to perform gradient-based optimization on multiple action proposals and select the best one.
+Moreover, we show that the reward model can either be learned from expert demonstrations or implemented using a zero-shot VLM, making GPC broadly applicable across diverse tasks.
 We evaluate GPC in both simulated and real-world, state-based and vision-based robotic manipulation tasks, demonstrating significant improvement over pure behavior cloning.
-Novelty.
-Among the three modules of GPC, the first one is not new and has become the common practice in robot policy learning [15, 52].
-The novelty of GPC lies in learning a predictive world model—crucially, using random exploration data—to strengthen the generative policy with online planning (as opposed to, e.g., offline RL [23, 28, 56, 17, 44]).
-Moreover, different from previous works leveraging generative models for world modeling, GPC learns a world model that is explicitly action-conditioned.
-In other words, the world model is “controllable” by the robot’s actions.
-We show such a controllable world model enables robust real-world robotic manipulation.
 Paper organization.
-We formalize the overall pipeline of GPC in §II and detail the approaches for state-based and vision-based world modeling in §III.
-We present experimental results in §IV and conclude in §VI.
-Review of related works and how GPC differs from them are described in §V.
-II Overview of Generative Predictive Control
+We formalize the overall pipeline of GPC in §2 and detail the approaches for state-based and vision-based world modeling in §3 and Appendix C.
+We present experimental results in §4 and conclude in §5.
+More results can be found in Appendix F and https://computationalrobotics.seas.harvard.edu/GPC/.
+Related works are reviewed in Appendix G.
+2 Overview of Generative Predictive Control
 Let $I_{t}$ be the information vector summarizing the state of the robot and the environment up to time step $t$.
@@
 Generative policy training.
-We ask humans to teleoperate robots to solve the tasks.
-The expert demonstrations are then divided into clips of data containing pairs of state information and action chunks, i.e., a dataset ${\cal D}_{\text{expert}}^{P}=\{I_{t}^{i}\leftrightarrow a_{t:t+T}^{i}\}_{i=1}^{N_{\text{expert}}}$ with number of samples $N_{\text{expert}}$.
-Policy learning becomes a supervised learning problem with input $I_{t}$ and output $a_{t:t+T}$.
-[^1] It is also common to include language instructions as input to the policy as in vision-language-action (VLA) models [32].
-GPC can generalize to VLAs in a straightforward way.
-We exclude language here to focus on motion planning.
-In GPC, we follow the diffusion policy learning framework [15].
-Specifically, we feed $I_{t}$ into a network that parametrizes the score function of the conditional distribution $p(a_{t:t+T}|I_{t})$, using which Gaussian noise vectors are gradually diffused to expert action chunks.
-The score function network is trained using denoising prediction [24].
-Policy learning leads to a policy network ${\cal P}(\cdot)$ that predicts action chunks $a_{t:t+T}$ given state information $I_{t}$.
-It is important to note that ${\cal P}(\cdot)$ is stochastic and effectively a sampler from the (approximate) conditional distribution $p(a_{t:t+T}|I_{t})$, meaning that given $I_{t}$, we can draw many possible action chunks $(a_{t:t+T}^{(1)},\dots,a_{t:t+T}^{(K)})$ from ${\cal P}(\cdot)$.
-We call each action chunk $a_{t:t+T}$ an action proposal.
+We ask humans to teleoperate robots to solve tasks, generating expert demonstrations that are segmented into clips of state-action pairs, forming a dataset ${\cal D}_{\text{expert}}^{P}=\{I_{t}^{i}\leftrightarrow a_{t:t+T}^{i}\}_{i=1}^{N_{\text{expert}}}$.
+Policy learning then reduces to supervised learning with input $I_{t}$ and output $a_{t:t+T}$.
+In GPC, we adopt the diffusion policy learning framework [7], where $I_{t}$ is fed into a network that parametrizes the score function of $p(a_{t:t+T}|I_{t})$.
+Gaussian noise vectors are gradually denoised into expert action chunks following [24].
+This yields a stochastic policy network ${\cal P}(\cdot)$ that samples action chunks $a_{t:t+T}$ given $I_{t}$.
+We refer to each sampled action chunk as an action proposal.
 Predictive world modeling.
-The fact that ${\cal P}(\cdot)$ is stochastic leads to a natural question: which one of the $K$ action proposals should be selected?
-The world model precisely aims to answer this question by predicting the future consequences of each action proposal.
-To build the world model, we need pairs of $(I_{t},a_{t:t+T})$ (current state information plus an action chunk) and $I_{t+1:t+T+1}$ (future state information).
-The expert demonstrations used for policy training already contain such pairs, i.e., a dataset ${\cal D}_{\text{expert}}^{W}=\{(I^{i}_{t},a^{i}_{t:t+T})\leftrightarrow I_{t+1:t+T+1}^{i}\}_{i=1}^{N_{\text{expert}}}$.
-Nevertheless, as we show in §IV, if we train the world model only using ${\cal D}_{\text{expert}}^{W}$, the predictive power of the world model is insufficient to guide action proposal selection.
-One can collect more expert demonstrations to enlarge the datasets, but it can be expensive.
-Instead, we posit that random exploration data are not only easier to collect, but also rich in dynamics.
-Therefore, we collect another exploration dataset ${\cal D}^{W}_{\text{explore}}=\{(I^{i}_{t},a^{i}_{t:t+T})\leftrightarrow I_{t+1:t+T+1}^{i}\}_{i=1}^{N_{\text{explore}}}$ where humans (or any other controllers) are asked to randomly perturb the system without solving any tasks.
-In control theory terminology, the goal of this is to perform “system identification” with “sufficient excitation” [37].
-We combine the two datasets ${\cal D}^{W}:={\cal D}^{W}_{\text{expert}}\cup{\cal D}^{W}_{\text{explore}}$ to train the world model.
-Details of the design of the world model architecture are deferred to §III.
-For now, the reader can assume we obtain a model ${\cal W}(\cdot)$ that predicts the future $I_{t+1:t+T+1}$ conditioned on the current state information and the action proposal $(I_{t},a_{t:t+T})$.
+The stochastic nature of ${\cal P}(\cdot)$ raises a key question: which of the many action proposals should be selected?
+The world model addresses this by predicting the future outcomes of each proposal.
+To train it, we use pairs of $(I_{t},a_{t:t+T})$ and $I_{t+1:t+T+1}$ from the expert dataset ${\cal D}_{\text{expert}}^{W}=\{(I^{i}_{t},a^{i}_{t:t+T})\leftrightarrow I_{t+1:t+T+1}^{i}\}_{i=1}^{N_{\text{expert}}}$, already available from policy training.
+However, as shown in §4, training solely on ${\cal D}_{\text{expert}}^{W}$ leads to limited predictive ability.
+While more expert data could help, it can be costly.
+Instead, we collect an additional exploration dataset ${\cal D}^{W}_{\text{explore}}=\{(I^{i}_{t},a^{i}_{t:t+T})\leftrightarrow I_{t+1:t+T+1}^{i}\}_{i=1}^{N_{\text{explore}}}$, where humans (or other controllers) randomly perturb the system without solving tasks.
+This approach, inspired from “system identification” with “sufficient excitation” in control theory [25], enriches the dynamics.
+We combine both datasets as ${\cal D}^{W}:={\cal D}^{W}_{\text{expert}}\cup{\cal D}^{W}_{\text{explore}}$ to train the world model.
+Architectural details are in §3; for now, we assume access to a model ${\cal W}(\cdot)$ that predicts $I_{t+1:t+T+1}$ from $(I_{t},a_{t:t+T})$.
 Online planning.
-We now formalize the two online planning algorithms, namely GPC-RANK and GPC-OPT, that combine the policy ${\cal P}(\cdot)$ and the world model ${\cal W}(\cdot)$ (cf.
-Fig.
-1).
-We assume a reward model ${\cal R}(\cdot):I_{t+1:t+T+1}\rightarrow r_{t}$ is available that predicts the reward given the future state information, e.g., a small neural network that is separately trained.
+We now formalize the two online planning algorithms, namely GPC-RANK and GPC-OPT, that combine the policy ${\cal P}(\cdot)$ and the world model ${\cal W}(\cdot)$.
+We assume a reward model ${\cal R}(\cdot):I_{t+1:t+T+1}\rightarrow r_{t}$ is available that predicts the reward given the future state information, e.g., a small neural network that is separately trained or foundation models (i.e., vision-language models (VLMs)) capable of zero-shot selection of the most promising future state.
+A trained reward predictor is appropriate in scenarios where the reward can be explicitly defined—particularly in numerical terms—and it is differentiable, enabling gradient-based action optimization.
+However, some tasks involve rewards that are difficult or even infeasible to specify.
+In such cases, VLMs can serve as surrogate reward predictors by directly selecting the most suitable action proposal, given the predicted outcomes and task description as input.
+This approach introduces greater flexibility to our GPC method, extending its applicability to a broader range of tasks.
   1. (i)
-    GPC-RANK is based on the simple intuition to pick the action proposal with highest reward. Formally, we sample $K$ action proposals from the policy
+    GPC-RANK is based on the intuition to pick the action proposal with highest reward. Formally, we sample $K$ action proposals from the policy
     $(a_{t:t+T}^{(1)},\dots,a_{t:t+T}^{(k)},\dots,a_{t:t+T}^{(K)})\sim{\cal P}(I_{t}),$ (1)
-    and output the best proposal by passing them through the world model ${\cal W}(\cdot)$ and computing respective rewards:
+    pass them through the world model ${\cal W}(\cdot)$, and select the one with highest reward:
     $\pi(I_{t})=a_{t:t+T}^{(k_{\star})},\quad k_{\star}\in\operatorname*{arg\,max}_{k=1,\dots,K}\ {\cal R}({\cal W}(I_{t},a_{t:t+T}^{(k)})),$ (2)
     where we used the notation $\pi(\cdot)$ to denote the online policy in contrast to the offline policy ${\cal P}(\cdot)$.
-The underlying assumption of GPC-RANK is that, due to distribution shifts, the optimal action chunk at test time may not have the highest likelihood in ${\cal P}(\cdot)$, and thus may need more samples to be discovered.
-The world model—assuming it is learned from a richer distribution because of random exploration—will pick up the optimal action chunk through (2).
-A nice property of GPC-RANK is that it can be easily parallelized.
+Due to distribution shifts, the optimal action chunk at test time may not correspond to the highest-likelihood sample from ${\cal P}(\cdot)$ and thus may require more samples to uncover.
+By training on a richer distribution via random exploration, the world model helps select the optimal chunk through (2).
+A key advantage of GPC-RANK is its easy parallelizability.
   2. (ii)
-    GPC-OPT takes a different perspective by directly solving the reward maximization problem given the world model:
+    GPC-OPT directly solves the reward maximization problem given the world model:
     $\max_{a_{t:t+T}}{\cal R}({\cal W}(I_{t},a_{t:t+T})),$ (3)
-    where the action chunk are viewed as decision variables.
-The difficulty is, however, problem (3) can be highly nonconvex due to the nonlinearity and nonsmoothness of ${\cal W}(\cdot)$ and ${\cal R}(\cdot)$.
-Fortunately, the pretrained policy ${\cal P}(\cdot)$ comes to rescue because it narrows down the search space by providing good initializations to solve (3), i.e., warmstarts.
-Formally, we sample an action chunk $\hat{a}_{t:t+T}$ from ${\cal P}(\cdot)$ and perform $M$ steps of gradient ascent starting from $\hat{a}_{t:t+T}$ (assuming ${\cal R}(\cdot)$ is differentiable):
-    $\hat{a}_{t:t+T}^{(0)}:=\hat{a}_{t:t+T}\sim{\cal P}(I_{t}),\text{ and for }\ell=1,\dots,M\text{ do:}$:
-    $\hat{a}_{t:t+T}^{(\ell)}=\hat{a}_{t:t+T}^{(\ell-1)}+\eta^{(\ell)}\nabla_{a_{t:t+T}}{\cal R}({\cal W}(I_{t},\hat{a}_{t:t+T}^{(\ell-1)})),$ (4)
-    where $\eta^{(\ell)}>0$ are step sizes. The online policy outputs the final optimized action chunk
+    treating the action chunk as decision variables.
+However, (3) is often highly nonconvex due to the complexity of ${\cal W}(\cdot)$ and ${\cal R}(\cdot)$.
+To address this, we leverage the pretrained ${\cal P}(\cdot)$ to warm-start optimization.
+We sample $\hat{a}_{t:t+T}$ from ${\cal P}(I_{t})$ and perform:
+    $\hat{a}_{t:t+T}^{(0)}:=\hat{a}_{t:t+T}\sim{\cal P}(I_{t}),\quad\hat{a}_{t:t+T}^{(\ell)}=\hat{a}_{t:t+T}^{(\ell-1)}+\eta^{(\ell)}\nabla_{a_{t:t+T}}{\cal R}({\cal W}(I_{t},\hat{a}_{t:t+T}^{(\ell-1)})),$ (4)
+    for $\ell=1,\dots,M$, with step sizes $\eta^{(\ell)}>0$. The final optimized action chunk is:
     $\pi(I_{t})=\hat{a}_{t:t+T}^{(M)}.$ (5)
-    Since the gradients in ((ii)) are readily available from automatic differentiation, gradient ascent can be carried out using off-the-shelf optimizers like ADAM.
-Readers familiar with the optimal control literature will recognize GPC-OPT is the classical single shooting algorithm [16].
-Compared to GPC-RANK which evaluates the world model $K$ times, GPC-OPT evaluates the world model $M$ times.
-The former can be parallelized but the latter cannot.
-GPC-RANK and GPC-OPT can be combined.
-We can sample $K$ action proposals from ${\cal P}(\cdot)$ and for each proposal perform reward maximization according to ((ii)), leading to $K$ optimized action chunks.
-We pick the optimized action chunk with highest reward.
-GPC-RANK+OPT can be viewed as solving the optimization (3) from multiple initializations [50], and requires $K\times M$ evaluations of the world model.
-We provide a succinct summary of GPC in Algorithm 1.
+    Gradients are obtained via automatic differentiation, and updates can be performed with optimizers like ADAM.
+GPC-OPT resembles the classical single shooting method [23].
+Unlike GPC-RANK, which evaluates the world model $K$ times in parallel, GPC-OPT evaluates it $M$ times sequentially.
+GPC-RANK and GPC-OPT can be combined by sampling $K$ action proposals from ${\cal P}(\cdot)$ and applying reward maximization (4) to each, yielding $K$ optimized action chunks.
+We select the one with the highest reward.
+GPC-RANK+OPT effectively solves (3) from multiple initializations [26], requiring $K\times M$ world model evaluations.
+A summary of GPC is provided in Appendix A, with an analysis of its connection to offline RL in Appendix E.
+The world model is crucial for ranking and optimizing action proposals.
+While early works such as deep visual foresight [27] explored learned world models, we next show that modern diffusion models now enable (near) physics-accurate visual predictions previously unattainable.
+3 World Model Learning
+The world model takes as input $(I_{t},a_{t:t+T})$ and predicts $I_{t+1:t+T+1}$.
+When $I_{t}$ represents a low-dimensional state, an MLP can be used to learn the dynamics, as demonstrated in Appendix C for the case of planar pushing.
+However, state-based world modeling relies on accurate state estimation, which is feasible in controlled lab environments with infrastructure such as AprilTags [28] or motion capture systems, but can be challenging in open or unstructured environments.
+Therefore, it is desirable to learn visual dynamics directly, i.e., to build a visual world model.
+Figure 2: Diffusion-based Visual World Modeling.
+Our visual world model consists of a sequence of recursive single-step image predictors.
+Each single-step image predictor is designed as a conditional diffusion process.
+Diffusion-based visual world modeling.
+Motivated by the success of using diffusion models for image generation, we design a visual world model based on conditional diffusion.
+Recall that the input to the world model is $I_{t}=o_{t-H:t}$ (the sequence of past images) and $a_{t:t+T}$, and the output is $I_{t+1:t+T+1}:=o_{t+1:t+T+1}$ (a sequence of future images).
+We design the visual world model as a recursive application of single-step image predictors, which reads
+$o_{t+1}=f_{\text{vision}}(o_{t-H:t},a_{t}),$ (6)
+where $f_{\text{vision}}$ is a conditional diffusion model.
+The unique property of diffusion is that $o_{t+1}$ is not generated in a single feedforward step, but rather through a sequence of denoising steps.
+In particular, let $o_{t+1}^{{\cal T}}$ be drawn from a white Gaussian noise distribution, then $f_{\text{vision}}$ proceeds as:
+$o_{t+1}^{\tau-1}=D_{\phi}(o_{t+1}^{\tau},\tau,o_{t-H:t},a_{t}),\tau={\cal T},\dots,1,$ (7)
+where ${\cal T}$ is the total number of denoising steps and the output of $f_{\text{vision}}$ is $o_{t+1}=o_{t+1}^{0}$.
+In (7), $\tau$ is the denoising step index and $D_{\phi}$—the “denoiser”—is a neural network with weights $\phi$.
+We use the same architecture for $D_{\phi}$ as [29], containing convolutions, action embedding, and a U-Net (Fig.
+2).
+$D_{\phi}$ is trained by adding random noises to the clean images and then predicting the noise.
+Two-phase training.
+To improve the accuracy and consistency of visual world modeling, we train it in two phases.
+In phase one, we train only the single-step image predictor (6), i.e., with supervision only from a single image $o_{t+1}$.
+In phase two, we recursively apply the single-step predictor (6) for $T$ times to obtain a sequence of future images $o_{t+1:t+T+1}$, and jointly supervise them with groundtruth images.
+We use observation horizon $H=4$ in the visual world modeling, and ${\cal T}=3$ diffusion steps.
+Remark 1 (Freeze the Noise).
+It should be emphasized that the single-step image predictor $f_{\text{vision}}$ in (6) is “stochastic” where the stochasticity comes from the randomness in the initial noise $o_{t+1}^{{\cal T}}$.
+In classical generative modeling, stochasticity is desired because we want the model to generate diverse images.
+However in our paper, we want to use the diffusion-based world model for control.
+That is, we want to understand the impact of the “action” on the future, rather than the impact of the random “noise” on the future.
+Therefore, we freeze the noise $o_{t+1}^{{\cal T}}$ to be zero at inference time, which makes the world model deterministic—outputting the most likely future predictions (because $o_{t+1}^{{\cal T}}=0$ attains the peak probability in the Gaussian distribution).
+In fact, without freezing the noise, GPC-OPT does not work because the gradient of the reward with respect to the actions (cf.
+(4)) would be “contaminated” by the noise and thus does not guide online planning.
+4 Experiments
+We evaluate GPC on (1) a state-based planar pushing task, (2) four vision-based simulation tasks, and (3) two real-world tasks.
+We found GPC consistently outperforms the behavior cloning baseline.
+Detailed descriptions of the tasks can be found in Appendix B.
+4.1 State-based Planar Pushing in Simulation
+We study the planar pushing task with the goal of pushing an object from an initial pose to a specified target pose, where the groundtruth pose of the object is available through a simulator.
+We firstly train a state-based diffusion policy ${\cal P}(\cdot)$ [7] from which we can sample multiple action proposals, as explained in §2.
+Then with both expert demonstrations and random exploration data, we utilize MLP networks to construct the state-based world model predicting the outcome states and define the reward with a registration loss between the target and predicted object states.
+Please refer to more training details in Appendix C.
+In Table 1, we present results for GPC, the pure behavior cloning baseline ($K=1$ and $M=0$), and Offline RL baseline (more details in Appendix E).
+Clearly, all GPC-RANK, GPC-OPT, and GPC-RANK+OPT outperform the baselines.
+Notably, GPC-RANK+OPT approaches the performance of planning with a groundtruth simulator.
+Offline RL GPC-RANK GPC-OPT GPC-RANK+OPT With GT simulator $K=100$ $K=5000$ $\begin{subarray}{c}K=1\\ M=0\end{subarray}$ $\begin{subarray}{c}K=100\\ M=0\end{subarray}$ $\begin{subarray}{c}K=1000\\ M=0\end{subarray}$ $\begin{subarray}{c}K=5000\\ M=0\end{subarray}$ $\begin{subarray}{c}K=1\\ M=30\end{subarray}$ $\begin{subarray}{c}K=10\\ M=30\end{subarray}$ $\begin{subarray}{c}K=20\\ M=30\end{subarray}$ $\begin{subarray}{c}K=30\\ M=20\end{subarray}$ Score $0.788$ $0.745$ $0.812$ 0.898 0.932 0.934 0.886 0.912 0.914 0.891 $0.952$
+Table 1: State-based Planar Pushing (Push-T) in Simulation.
+Score is measured by the IoU metric averaged over $100$ evaluation seeds.
+$K$ is the number of action proposals sampled from ${\cal P}(\cdot)$ while $M$ is the number of gradient steps performed to maximize reward.
+4.2 Vision-based Tasks in Simulation
+Figure 3: World model prediction in GPC for simulation tasks.
+Push-T Triangle Drawing Block Stacking Cube & Sphere Swapping Reward Learned Learned VLM Learned VLM VLM Behavior Cloning $0.642$ $0.724$ $0.724$ $0.912$ $0.912$ 0.68 GPC-RANK 0.739 0.775 0.755 0.988 0.987 0.73
+Table 2: Scores for four vision-based tasks in simulation.
+Figure 4: Comparison of different visual world modeling on two examples. (Highlights for comparison are boxed in blue).
+We test on four vision-based simulation tasks named as Push-T, Triangle Drawing, Block Stacking, and Cube & Sphere Swapping.
+We train a vision-based diffusion policy ${\cal P}(\cdot)$ [7] using ResNet18 and UNet.
+For the world model, as in §3, we use a diffusion model to build a single-step image predictor and recursively apply it $T$ times to form the world model ${\cal W}(\cdot)$.
+Same as in §4.1, the training data for world model includes random explorations.
+Please see more training details in Appendix D.
+Reward.
+Different from the state-based planar pushing, where future states are structured, the image-based predictions make it difficult to directly compute a reward signal.
+Therefore, we propose two ways to implement the reward function.
+(1) When a numerical reward can be defined, such as the registration loss between the T-block and the target frame in Push-T, or the distance between two cubes in Block Stacking, we train a reward predictor with a ResNet18 backbone followed by an MLP, mapping images to scalar rewards.
+The learned reward model $\mathcal{R}(\cdot)$ is differentiable and can be used for online planning.
+(2) For tasks where numerical rewards are tedious or infeasible to define, we propose leveraging VLM (ChatGPT-4o [30]) to select among action proposals in a zero-shot manner.
+After predicting outcomes with the world model, we prompt the VLM with the predicted images and the task goal, allowing it to choose the best action proposal.
+Quality of visual world modeling.
+Fig.
+3 illustrates the predicted images generated by our world model, demonstrating both visually compelling results and accurate predictions of object interactions and movements.
+Then we study whether our visual world modeling technique is better than some baselines.
+Particularly, we chose two baselines, one from the pioneering deep visual foresight [27], and the other from the more recent AVDC paper [15].
+The world model in deep visual foresight leverages CNN and LSTM for future prediction, and we follow the Github implementation.
+[^1] https://github.com/Xiaohui9607/physical_interaction_video_prediction_pytorch AVDC is originally designed for video diffusion conditioned on language, and we modify its official implementation [^2] https://github.com/flow-diffusion/AVDC to condition on robot actions.
+Notably, AVDC is a multiple-step world model as it simultaneously diffuses future images together, unlike our recursive structure.
+Fig.
+4 compares the quality of future predictions of three methods (ours, deep visual foresight, and AVDC) against the groundtruth in the Push-T task.
+We see that diffusion-based visual world modeling (ours and AVDC) clearly dominates the classical deep visual foresight.
+Between ours and AVDC, we can observe that the future predictions from our method are better aligned with the groundtruth.
+Planning performance.
+We present the results of GPC-RANK and pure behavior cloning baseline in Table 2.
+With either a learned reward predictor ${\cal R}(\cdot)$ to pick from $100$ action proposals or VLM with $10$ candidates, GPC-RANK can significantly improve the pure behavior cloning baseline for all four vision-based tasks.
+Fig.
+5 shows evaluation trajectories comparing GPC-RANK and baseline.
+Figure 5: Evaluation Trajectories for Triangle Drawing and Cube & Sphere Swap.
+Baseline Methods Method Offline RL ($K=100,M=0$) Behavior Cloning ($K=1,M=0$) CEM ($K=100,M=0$) CEM ($K=200,M=0$) Gradient Ascent ($K=1,M=25$) Score $0.620$ $0.642$ $0.174$ $0.202$ $0.131$
+GPC Method GPC-RANK ($K=50,M=0$) GPC-RANK ($K=100,M=0$) GPC-OPT ($K=1,M=25$) GPC-RANK+OPT ($K=10,M=25$) Score 0.698 0.739 0.791 0.882
+Table 3: Ablation Study for Vision-based Planar Pushing (Push-T) in Simulation.
+Score is measured by the IoU metric averaged over 100 evaluation seeds.
+$K$ is the number of action proposals and $M$ is the number of gradient steps.
+Detailed comparison in planar pushing. We study the impact of $K$ and $M$ on the performance of GPC, its comparison with additional baselines, and the importance of random exploration.
+Figure 6: Importance of random exploration in world model learning (vision-based push-T)
+In Table 3, we study how different variations of $K$ and $M$ affect the performance of GPC.
+Since image operations require significantly more memory and computation than state-based inputs, we limit $K$ to the scale of hundreds in vision-based experiments.
+The results show that: (a) GPC-RANK improves performance by approximately $10\%$ over the diffusion policy baseline ($K=1$, $M=0$); (b) GPC-OPT yields a $15\%$ improvement; and (c) GPC-RANK+OPT achieves up to a $25\%$ gain.
+Table 3 also includes comparisons with four baselines: Offline RL [31], pure behavior cloning [7], the cross entropy method (CEM) [32], and pure gradient ascent [32].
+These results indicate that neither behavior cloning, offline RL, nor standalone planning methods are sufficient to achieve high performance in complex vision-based tasks.
+In contrast, GPC, which unifies generative and predictive control, consistently outperforms these baselines.
+In Fig.
+6, we compare the performance of GPC-RANK and GPC-OPT for Push-T when trained with or without random exploration.
+The performance is increased by $10\%$ approximately when introducing random exploration data, which demonstrates its importance in achieving an accurate world model.
+Similar results for state-based task are presented in Appendix C.
+4.3 Real-world Vision-based Tasks
+We then test GPC on real-world tasks, involving push-T (and its variations with object collisions) and clothes folding.
+For the push-T tasks, rewards are computed using a learned reward predictor composed of a ResNet and MLP, trained with registration losses derived from expert demonstrations’ state information using AprilTags [28].
+However for cloth folding, where states are harder to define, we design a progressing reward: starting at 00 for spread clothes and increasing with each successful grasp-and-move operation.
+Normalized progressing rewards for expert demonstrations are used to train the reward predictor for clothes folding.
+For plain push-T, while the baseline model succeeds 5 out of 10 times, both GPC-RANK and GPC-OPT succeeds 7 out of 10 times.
+In a more challenging version of push-T with obstacles (letters "A" and "R") placed on the table, out of 5 test trials, the baseline succeeds 2 times, GPC-RANK succeeds 3 times, and GPC-OPT succeeds 4 times.
+For the clothes folding task, while the baseline model succeeds 3 out of 10 times, GPC-RANK succeeds 7 out of 10 times.
+Fig.
+7 and Fig.
+8 show the trajectories for these tests.
+Despite the challenging nature of these real-world experiments—where dynamics are more complex due to collisions and tasks like cloth folding involve non-rigid bodies and require multiple steps—GPC still operates effectively.
+More trajectories and the quality of world model prediction for real-world tests are shown in Appendix F and https://computationalrobotics.seas.harvard.edu/GPC/.
+Figure 7: Real-world tests for push-T with collision.
+In every test, top row shows trajectories of baseline model ($K=1,M=0$), middle row shows trajectories of GPC-RANK ($K=10,M=0$), and last row shows trajectories of GPC-OPT ($K=0,M=25$).
+Figure 8: Real-world tests for clothes folding. Top row shows trajectories of baseline model ($K=1,M=0$), second row shows trajectories of GPC-RANK ($K=10,M=0$).
+5 Conclusion
+We presented generative predictive control (GPC), a learning-based control paradigm that integrates a generative policy with a predictive world model for online planning.
+While GPC shares strong conceptual ties with classical predictive control, namely identifying a dynamics model followed by online optimization, it introduces several novel insights.
+The new insights offered by GPC are that (a) modern conditional diffusion models provides a powerful tool to learn controllable action-conditioned visual world models; (b)VLM based method can be used to pick from imagined outcomes of visual world models, as an alternative to learned reward predictors; and (c) a generative policy trained from behavior cloning provides good warmstarts for online optimization that helps mitigate the challenges posed by nonconvex reward landscapes in online optimization.
+Through extensive experiments in both simulated and real-world settings, we demonstrated the robustness and versatility of GPC in both state-based and vision-based robotic manipulation.
+6 Limitation
+Two issues are worth highlighting.
+The first issue is about hallucination.
+One of the important issues of generative modeling is hallucination, and our visual world model is no exception.
+In particular, there are several cases where the world model predicts images of the future that violate rigid-body physics.
+For example, the objects T, A, and R may penetrate each other, as we show in Fig.
+9.
+Interestingly, even though the predictions can violate physics, using such predictions for planning still leads to GPC’s superior performance.
+We believe learning a visual world model that strictly satisfies the laws of physics remains a challenging task, and strategies such as data scaling and baking in physics priors may help solve this challenge.
+For example, it may be possible to leverage the so-called world foundation models [33]—trained on Internet-scale videos—and finetune them for specific robotic tasks.
+The second issue is associated with the computational complexity of using an expensive diffusion-based world model in online planning.
+It is worth noting that, although GPC dominates the baseline in terms of success rates, it is generally slower than the baseline due to predicting the future using the world model.
+A promising solution is to leverage diffusion distillation diffusion [34, 35], and faster diffusion solvers [36] for faster world model prediction.
+Figure 9: Limitations of world model prediction.
+Appendix A GPC Pseudocode
 Input: Expert action demonstrations ${\cal D}^{P}_{\text{expert}}=\{I_{t}^{i}\leftrightarrow a^{i}_{t:t+T}\}_{i=1}^{N_{\text{expert}}}$; Expert trajectories and random exploration trajectories ${\cal D}^{W}_{\text{expert}}\cup{\cal D}^{W}_{\text{explore}}=\{(I_{t}^{i},a^{i}_{t:t+T})\leftrightarrow I^{i}_{t+1:t+T+1}\}_{i=1}^{N_{\text{expert}}+N_{\text{explore}}}$; Reward model ${\cal R}$; Positive integers $K$, $M$
@@
 4
-5// Predictive world modelingIII)
+5// Predictive world modeling3)
 […]
@@
 Algorithm 1 Generative Predictive Control (GPC)
-Remark 1 (Connection with Offline RL).
-It is worth noting that the composition of the reward model ${\cal R}(\cdot)$ and the world model ${\cal W}(\cdot)$ can be identified as the so-called Q-value function in offline RL, as they predict the reward of an action chunk $a_{t:t+T}$ given the current information $I_{t}$.
-However, there are three subtle differences between GPC and offline RL.
-(a) GPC learns an explicit dynamics model instead of just approximating the Q-value function (this is similar to model-based offline RL).
-(b) The composition of ${\cal R}(\cdot)$ and ${\cal W}(\cdot)$ provides a Q-value function for “action chunks” while the usual Q-value function involves a single action.
-(c) GPC learns the dynamics model using diverse trajectories (without rewards), instead of only expert demonstrations.
-In §IV, we compare GPC with an offline RL method adapted from single action to action chunk and show that GPC outperforms offline RL.
-The above overview of GPC highlights the importance of the world model in both ranking and optimizing action proposals.
-Indeed, the concept of a learned world model dates back to early pioneering works on deep visual foresight [20].
-In the next section, we will show that modern diffusion models are the key enabler for learning (near) physics-accurate visual world models that were not possible before.
-Figure 2: Illustration of planar pushing.
-Figure 3: Diffusion-based Visual World Modeling.
-Our visual world model consists of a sequence of recursive single-step image predictors.
-Each single-step image predictor is designed as a conditional diffusion process.
-III World Model Learning
-Recall that the world model takes as input $(I_{t},a_{t:t+T})$ and outputs $I_{t+1:t+T+1}$.
-We will briefly present the world model design when $I_{t}=x_{t}$ contains low-dimensional states (§III-A) and focus on a diffusion-based world model when $I_{t}=o_{t-H:t}$ contains high-dimensional images (§III-B).
-Planar pushing.
-Before getting started, since we will focus on planar pushing in our experiments §IV, we briefly introduce the problem setup.
-As shown in Fig.
-2, the planar pushing task involves pushing an object (e.g., a T block) to align with the target using a pusher.
+Appendix B GPC Task Descriptions
+In this section, we describe the 4 simulation tasks and 2 real-world tasks including the variations.
+B.1 Simulation tasks
+  • Push-T.
+This is a planar pushing task involving pushing an object (e.g., a T block) to align with the target using a pusher.
+We use the same simulator as in [7] with a T-shape object.
 We assume the target position is fixed and known, and hence we do not need to predict it.
-The state of the object is fully described by its 2D coordinates $p=(p_{x},p_{y})$ and 1D angle $\theta$.
+The low-dimensional state of the object is fully described by its 2D coordinates $p=(p_{x},p_{y})$ and 1D angle $\theta$.
 The action space is $a\in\mathbb{R}^{2}$ containing pushing velocities in the $x$ and $y$ directions.
-We choose the planar pushing task for two reasons.
-First, it is easy to understand and simple to visualize.
-Second, it is also a long-standing difficult task because it is contact-rich.
+Push-T is a long-standing difficult task because it is contact-rich.
 Depending on whether the pusher makes or breaks contact and whether the contact is sticking or sliding, the dynamics of the task involves multiple different modes and requires intricate physics reasoning [37].
-This makes planar pushing a both elementary and challenging task for world model learning.
-III-A State-based World Modeling
-For state-based world modeling, with $x=(p,\theta)$ being the state of the object, we learn a single-step dynamics model:
-$x_{t+1}=f_{\text{state}}(x_{t},a_{t}).$ (6)
-As pointed out by [64], when rotations are involved in machine learning, it is crucial to choose a rotation representation that is continuous.
-For example, the angle representation $\theta$ is discontinuous due to the “wrap $2\pi$” issue.
-Therefore, we augment the state as
-$x=(p,\cos(\theta),\sin(\theta))$ (7)
-for both the input and output of $f_{\text{state}}$ in (6) (we normalize the output of $f_{\text{state}}$ to satisfy $\sin^{2}(\cdot)+\cos^{2}(\cdot)=1$).
-We also found it beneficial to have two separate networks, one predicting the position and the other predicting the rotation.
-In our implementation, for each of the network, we used a six-layer MLP with hidden dimension $128$.
-During testing, we need to predict $T$ steps into the future given a sequence of actions $a_{t:t+T}$.
-We do so by recursively applying the single-step predictor (6).
-As we will show in §IV, this simple MLP-based single-step predictor is sufficient for robust online planning.
-III-B Vision-based World Modeling
-State-based world modeling requires accurate state estimation, which is possible in a lab environment with infrastructure such as AprilTags [46] or motion capture but can be challenging in the open environment.
-Therefore, it is desired to directly learn visual dynamics, i.e., a visual world model.
-Diffusion-based visual world modeling.
-Motivated by the success of using diffusion models for image generation, we design a visual world model based on conditional diffusion.
-Recall that the input to the world model is $I_{t}=o_{t-H:t}$ (the sequence of past images) and $a_{t:t+T}$, and the output is $I_{t+1:t+T+1}:=o_{t+1:t+T+1}$ (a sequence of future images).
-Similar as the state-based world modeling, we design the visual world model as a recursive application of single-step image predictors.
-We denote the single-step predictor as
-$o_{t+1}=f_{\text{vision}}(o_{t-H:t},a_{t}),$ (8)
-where $f_{\text{vision}}$ is a conditional diffusion model.
-The unique property of diffusion is that $o_{t+1}$ is not generated in a single feedforward step, but rather through a sequence of denoising steps.
-In particular, let $o_{t+1}^{{\cal T}}$ be drawn from a white Gaussian noise distribution, then $f_{\text{vision}}$ proceeds as:
-$o_{t+1}^{\tau-1}=D_{\phi}(o_{t+1}^{\tau},\tau,o_{t-H:t},a_{t}),\tau={\cal T},\dots,1,$ (9)
-where ${\cal T}$ is the total number of denoising steps and the output of $f_{\text{vision}}$ is $o_{t+1}=o_{t+1}^{0}$.
-In (9), $\tau$ is the denoising step index and $D_{\phi}$—the “denoiser”—is a neural network with weights $\phi$.
-We use the same architecture for $D_{\phi}$ as in [5], which contains convolutions, action embedding, and a U-Net, as illustrated in Fig.
-3.
-The denoiser $D_{\phi}$ is trained by adding random noises to the clean images and then predicting the noise.
-Two-phase training.
-To improve the accuracy and consistency of visual world modeling, we train it in two phases.
-In phase one, we train only the single-step image predictor (8), i.e., with supervision only on a single image $o_{t+1}$.
-In phase two, we recursively apply the single-step predictor (8) for $T$ times to obtain a sequence of future images $o_{t+1:t+T+1}$, and jointly supervise them with groundtruth images.
-We use observation horizon $H=4$ in the visual world modeling, and ${\cal T}=3$ diffusion steps in $f_{\text{vision}}$.
-Remark 2 (Freeze the Noise).
-It should be emphasized that the single-step image predictor $f_{\text{vision}}$ in (8) is “stochastic” where the stochasticity comes from the randomness in the initial noise $o_{t+1}^{{\cal T}}$.
-In classical generative modeling, stochasticity is desired because we want the model to generate diverse images.
-However in our paper, we want to use the diffusion-based world model for control.
-That is, we want to understand the impact of the “action” on the future, rather than the impact of the random “noise” on the future.
-Therefore, we freeze the noise $o_{t+1}^{{\cal T}}$ to be zero at inference time, which makes the world model deterministic—outputing the most likely future predictions (because $o_{t+1}^{{\cal T}}=0$ attains the peak probability in the Gaussian distribution).
-In fact, without freezing the noise, GPC-OPT does not work because the gradient of the reward with respect to the actions (cf.
-((ii))) would be “contaminated” by the noise and thus does not guide online planning.
-IV Experiments
-In this section, we focus on evaluation of the GPC framework on planar pushing and comparison to baseline behavior cloning.
-As mentioned in §III, we choose planar pushing because it is an elementary but also quite challenging task due to the complicated contact and collision dynamics involved.
-IV-A State-based Planar Pushing in Simulation
+The performance score is evaluated by the IoU value between block and target.
+  • Triangle Drawing.
+This is a simulation task built with ManiSkill simulator [38].
+The robotic hand is holding a pen to trace the triangle pattern on the paper.
+This is a task having high requirement about vision since it needs to trace triangle appearing in different location and orientation.
+The performance score is evaluated by the collapsed pixels between drawing and pattern.
+  • Block Stacking.
+This [39] is a classical manipulation task, aiming to stack one block onto another target block.
+This involves moving the block exactly above the target and and putting the block down.
+It is also built with the ManiSkill simulator [38].
+The performance score is evaluated by the distance between two blocks.
+  • Cube & Sphere Swap.
+This is also a simulation task built with ManiSkill simulator [38].
+Initially, there is a cube in the box and a sphere outside of the box.
+It requires the robot arm to move the cube out of the box and then move the sphere into the box.
+This task is challenging because it requires a logical operation sequence.
+Moreover, it is a multi-object tasks, which is challenging for the world-model prediction.
+The performance score is evaluated by completion status: $1.0$ is fully finished, $0.5$ is the half completion (only cube is moved out), $0.0$ is not done.
+B.2 Real-world tasks
+  • Plain Push-T.
+Since planar pushing is a challenging and elementary task, we also include the push-T task and its variations in real-world experiments.
+The pusher is attached to the end-effector of a Franka Panda robotic arm to push a T block into the target position.
+An overhead Intel RealSense camera is used to provide image observations to the pusher.
+To train the reward model, we need to obtain block information as training data so we attached three AprilTags [28] to the T-block, which enables accurate pose estimation from image frames even in the case of occlusion.
+To collect real-world expert demonstrations, we follow Chi et al.
+[7] and use a SpaceMouse to teleoperate the robotic arm to push the T block.
+A and R blocks are the possible objects that T may collide into during pushing.
+  • Push-T with obstacle A.
+To make task more challenging, we put an object of shape “A” on top of the target position so that when pushing T, it will inevitably collide with object A.
+The world model here not only needs to learn the dynamics for pushing T block, but also capture the dynamics about interactions between T-block and A-block.
+  • Push-T with obstacle A and R.
+Similarly, we put an object of shape “A” and “R” on top of the target position so that when pushing T, it will inevitably collide with object A and R.
+Here the dynamics is even more complex.
+  • Push-T in unseen scenarios.
+Here we use the policy and world model trained with push-T, push-T colliding A, push-T colliding A & R expert demos.
+But we perform the test to push T when an object of letter “R” is on top of the target.
+This situation has never been seen in any of the training data.
+  • Clothes Folding.
+We put a piece of clothes in random position and orientation on the table and ask the Franka hand to fold the clothes.
+It involves three steps: fold the left arm, fold the right arm, and fold the cap.
+Moreover, this is a task about operating on non-rigid body, making it quite challenging for both policy and world model.
+Appendix C State-based Planar Pushing
 We first study the setup where the groundtruth state of the object (i.e., $(p_{x},p_{y},\theta)$ is available through a simulator. We use the same simulator as in [7] with a T-shape object.
 Policy.
-We first train a state-based diffusion policy ${\cal P}(\cdot)$ [15] from which we can draw action proposals, as explained in §II.
+We first train a state-based diffusion policy ${\cal P}(\cdot)$ [7] from which we can draw action proposals, as explained in §2.
 We use $500$ expert demonstrations to train the policy network for $300$ epochs with AdamW optimizer of learning rate $10^{-4}$ and weight decay $10^{-6}$.
@@
 Then we need the world model to predict the future consequences of each action proposal.
-As in §III-A, we use MLP networks to build a single-step state predictor and recursively apply it $T$ times to form the world model ${\cal W}(\cdot)$.
-As introduced in §II, to ensure the dynamics of the world model is learned from a richer distribution than the expert demonstrations, the training dataset of the world model is expanded with random exploration demonstrations which attempt to record more pusher-object interactions at different contact points.
+For state-based world modeling, since we focus on planar pushing tasks here, we define with $x=(p,\theta)$ being the state of the object, then we learn a single-step dynamics model:
+$x_{t+1}=f_{\text{state}}(x_{t},a_{t}).$ (A1)
+As pointed out by [40], when rotations are involved in machine learning, it is crucial to choose a rotation representation that is continuous.
+For example, the angle representation $\theta$ is discontinuous due to the “wrap $2\pi$” issue.
+Therefore, we augment the state as
+$x=(p,\cos(\theta),\sin(\theta))$ (A2)
+for both the input and output of $f_{\text{state}}$ in (A1) (we normalize the output of $f_{\text{state}}$ to satisfy $\sin^{2}(\cdot)+\cos^{2}(\cdot)=1$).
+We also found it beneficial to have two separate networks, one predicting the position and the other predicting the rotation.
+In our implementation, for each of the network, we used a six-layer MLP with hidden dimension $128$.
+During testing, we need to predict $T$ steps into the future given a sequence of actions $a_{t:t+T}$.
+We do so by recursively applying the single-step predictor (A1).
+As we will show in §4, this simple MLP-based single-step predictor is sufficient for robust online planning.
+As introduced in §2, to ensure the dynamics of the world model is learned from a richer distribution than the expert demonstrations, the training dataset of the world model is expanded with random exploration demonstrations which attempt to record more pusher-object interactions at different contact points.
 To be specific, we let human operators (who are different from the operator of the expert demonstrations) collect $7$ long random play demonstrations and 1 special long demonstration that randomly collides the object with the boundary wall.
@@
 Results.
-In Table I, we study how different variations of $K$ and $M$ affect the performance.
+In Table 1, we study how different variations of $K$ and $M$ affect the performance.
 When $K=1$ and $M=0$, this is the baseline state-based diffusion policy trained from behavior cloning [7].
@@
 We also tried larger $K$, such as $1000$ and $5000$, and the performance further increases to $0.934$.
-Note that in the last column of Table I, we provided a “performance upper bound” which is obtained by using the groundtruth simulator (instead of the learned world model) to rank action proposals at $K=5000$.
+Note that in the last column of Table 1, we provided a “performance upper bound” which is obtained by using the groundtruth simulator (instead of the learned world model) to rank action proposals at $K=5000$.
 For GPC-OPT, if we sample one action proposal and perform gradient ascent from it (e.g., doing $30$ gradient steps), the score can be increased to $0.886$, which is also quite significant compared to the baseline.
@@
 The improvement remains substantial, indicating the robustness of GPC and shows both sampling multiple proposals and gradient-based optimization consistently improve the performance.
-In the first two columns of Table I, we present the results for another baseline of using Offline RL to rank and select actions.
-The Offline RL method follows that of [44] and is described in Appendix -A.
+In the first two columns of Table 1, we present the results for another baseline of using Offline RL to rank and select actions.
+The Offline RL method follows that of [31] and is described in Appendix E.
 Crucially, the Offline RL method only uses the expert demonstrations and does not leverage the random exploration data (because the random exploration data do not have labeled rewards).
@@
 A plausible reason for this is that the Offline RL method does not explicitly seek to learn the dynamics from the random explorations and hence the learned value function may be misleading for ranking action proposals.
-Finally, in Table II, we present the results of using a world model trained without random exploration data.
+Finally, in Table 4, we present the results of using a world model trained without random exploration data.
 We observe that, when the world model is only trained from expert demonstrations, the performance of both GPC-RANK and GPC-OPT gets worse.
 This demonstrates the importance of using random exploration data to learn an accurate world model.
-Offline RL GPC-RANK GPC-OPT GPC-RANK+OPT With GT simulator $K=100$ $K=5000$ $\begin{subarray}{c}K=1\\ M=0\end{subarray}$ $\begin{subarray}{c}K=100\\ M=0\end{subarray}$ $\begin{subarray}{c}K=1000\\ M=0\end{subarray}$ $\begin{subarray}{c}K=5000\\ M=0\end{subarray}$ $\begin{subarray}{c}K=1\\ M=30\end{subarray}$ $\begin{subarray}{c}K=10\\ M=30\end{subarray}$ $\begin{subarray}{c}K=20\\ M=30\end{subarray}$ $\begin{subarray}{c}K=30\\ M=20\end{subarray}$ Score $0.788$ $0.745$ $0.812$ $0.898$ $0.932$ $0.934$ $0.886$ $0.912$ $0.914$ $0.891$ $0.952$
-TABLE I: State-based Planar Pushing (Push-T) in Simulation.
-Score is measured by the IoU metric averaged over $100$ evaluation seeds.
-$K$ is the number of action proposals sampled from ${\cal P}(\cdot)$ while $M$ is the number of gradient steps performed to maximize reward.
 GPC-RANK GPC-OPT $K=100,M=0$ $K=1000,M=0$ $K=5000,M=0$ $K=1,M=30$ Score (using world model trained without random exploration) $0.832$ $0.838$ $0.816$ $0.752$ Score (using world model trained with random exploration) $0.898$ $0.932$ $0.934$ $0.886$
-TABLE II: The importance of random exploration in world model learning (state-based Push-T).
-IV-B Vision-based Planar Pushing in Simulation
-Figure 4: Comparison of different visual world modeling techniques on two examples.
-We then proceed to vision-based planar pushing with the same T-shape object.
+Table 4: The importance of random exploration in world model learning (state-based Push-T).
+Appendix D Vision-based Simulation Tasks
+We test on four vision-based simulation tasks, named as Push-T, Triangle Drawing, Block Stacking, and Cube & Sphere Swapping.
 Policy.
 We train a vision-based diffusion policy ${\cal P}(\cdot)$ [7] using ResNet18 as the vision encoder and UNet as the action diffusion network.
-We used $500$ expert demonstrations to train the policy for $300$ epochs with AdamW optimizer of learning rate $10^{-4}$ and weight decay $10^{-6}$.
+We train the policy for $300$ epochs with AdamW optimizer of learning rate $10^{-4}$ and weight decay $10^{-6}$.
 We use observation horizon of 2, prediction horizon of 16, and action horizon of 9.
 World model.
-As in §III-B, we use a diffusion model to build a single-step image predictor and recursively apply it $T$ times to form the world model ${\cal W}(\cdot)$.
-Same as in §IV-A, the training data includes random explorations.
+As in §3, we use a diffusion model to build a single-step image predictor and recursively apply it $T$ times to form the world model ${\cal W}(\cdot)$.
+Same as in §4.1, the training data includes random explorations.
 Reward.
 Different from the state-based planar pushing, since the future predictions are images, we cannot directly compute a reward signal.
-Therefore, we first train a pose prediction network with ResNet18 plus MLP, and then compute the registration loss using the predicted object pose.
+Therefore, for push-T, we first train a pose prediction network with ResNet18 plus MLP, and then compute the registration loss using the predicted object pose.
+For triangle drawing, we obtain the reward for expert demonstrations, defined as the number of pixels drawn on the pattern.
+For block stacking, we obtain the reward for expert demonstrations defined as the distance between two blocks.
+Then we train a reward predictor to predict the reward from any input images.
 Note that this learned reward model ${\cal R}(\cdot)$ is still differentiable.
-Online planning: With the vision-based policy, the visual world model, and the learned reward model, we follow the same online planning protocol described in §IV-A.
-Quality of visual world modeling.
-Before presenting the results of GPC for online planning, we first study whether our visual world modeling technique is better than some baselines.
-Particularly, we chose two baselines, one from the pioneering deep visual foresight [20], and the other from the more recent AVDC paper [19].
-The world model in deep visual foresight leverages CNN and LSTM for future prediction, and we follow the Github implementation.
-[^3] https://github.com/Xiaohui9607/physical_interaction_video_prediction_pytorch AVDC is originally designed for video diffusion conditioned on language, and we modify its official implementation [^4] https://github.com/flow-diffusion/AVDC to condition on robot actions.
-Notably, AVDC is a multiple-step world model as it simultaneously diffuses future images together, unlike our recursive structure.
-Fig.
-4 compares the quality of future predictions of three methods (ours, deep visual foresight, and AVDC) against the groundtruth in two examples.
-We see that diffusion-based visual world modeling (ours and AVDC) clearly dominates the classical deep visual foresight.
-Between ours and AVDC, we can observe that the future predictions from our method are better aligned with the groundtruth.
-Planning performance. Same as §IV-A, we study the impact of $K$ and $M$ on the performance of GPC, its comparison with Offline RL, and the importance of random exploration.
-In Table III, we study how different variations of $K$ and $M$ affect the performance of GPC.
-We can see that (a) GPC-RANK increases around $15\%$ from the diffusion policy baseline (when $K=1,M=0$); (b) GPC-OPT increases around $25\%$ from the diffusion policy baseline; and (c) GPC-RANK+OPT increases up to $40\%$ from the baseline.
-Contrasting the vision-based results with the state-based results, we observe that GPC-OPT appears to have a larger advantage in vision-based push-T.
-This indeed makes sense because the baseline vision-based diffusion policy only has a score of $0.642$, leaving a lot of room for optimization to find better actions.
-In the first column of Table III, we show the results of Offline RL. Unfortunately, Offline RL still does not perform as well as even the baseline vision-based diffusion policy.
-In Table IV, we compare the performance of GPC-RANK and GPC-OPT when trained with or without random exploration.
-Consistent with our observations in IV-A, without training on random exploration, the performance of online planning significantly decreases.
-Offline RL GPC-RANK GPC-OPT GPC-RANK+OPT $K=100$ $K=1,M=0$ $K=50,M=0$ $K=100,M=0$ $K=200,M=0$ $K=1,M=25$ $K=5,M=10$ $K=10,M=25$ Score $0.620$ $0.642$ $0.698$ $0.739$ $0.719$ $0.791$ 0.824 0.882
-TABLE III: Vision-based Planar Pushing (Push-T) in Simulation.
-Score is measured by the IoU metric averaged over $100$ evaluation seeds.
-$K$ is the number of action proposals and $M$ is the number of gradient steps.
-GPC-RANK GPC-OPT $K=100,M=0$ $K=1,M=25$ Score (using world model trained without random exploration) $0.647$ 0.652 Score (using world model trained with random exploration) $0.739$ $0.791$
-TABLE IV: The importance of random exploration in world model learning (vision-based push-T).
-IV-C Real-world Vision-based Planar Pushing
-We then test GPC in real-world planar pushing. We not only show GPC is effective in “plain push-T”, but also show it can work for push-T with collision objects.
-Figure 5: Real-world setup.
-Setup.
-Fig.
-5 depicts our real-world setup for vision-based planar pushing.
-The pusher is attached to the end-effector of a Franka Panda robotic arm to push a T block into the target position.
-An overhead Intel RealSense camera is used to provide image observations to the pusher.
-To train the reward model, we need to obtain block information as training data so we attached three AprilTags [46] to the T-block, which enables accurate pose estimation from image frames even in the case of occlusion.
-To collect real-world expert demonstrations, we follow Chi et al.
-[15] and use a SpaceMouse to teleoperate the robotic arm to push the T block.
-A and R blocks are the objects that T may collide into during pushing.
-IV-C1 Plain push-T
-We first test on plain push-T (i.e., without collision with other objects).
-We train the policy with $100$ expert push-T demonstrations using the architecture mentioned in §IV-B.
-Because of the slower moving speed of real-world pusher, we use a larger prediction horizon $T=32$.
-Same as §IV-B, we train the world model using all expert demonstrations and a few random explorations.
-For the reward model, we rely on AprilTag [46] to obtain the ground-truth object poses and use them to train the reward model ${\cal R}(\cdot)$.
-To be clear, our GPC pipeline in real world is still purely vision-based and the AprilTags are only used to obtain training data for the reward model (training the pose predictor).
-We follow the same online planning protocol described in §IV-A.
-We test GPC-RANK and GPC-OPT on a set of 10 push-T tasks where the T block is initialized at positions not seen during training.
-Performance.
-While the baseline model succeeds 5 out of 10 times, GPC-RANK succeeds 7 out of 10 times and GPC-OPT succeeds 7 out of 10 times.
-Fig.
-10(a) shows the trajectories for one test.
-More testing trajectories are shown in Appendix -B and supplementary videos.
+Besides, we propose VLM can be used to decide which action proposal to pick in a zero-shot way.
+We utilize ChatGPT to pick the best from predicted consequences for $10$ actions proposals.
+Online planning: With the vision-based policy, the visual world model, and the learned reward model, we follow the same online planning protocol described in §C.
+Appendix E Offline RL
+It is worth noting that the composition of the reward model ${\cal R}(\cdot)$ and the world model ${\cal W}(\cdot)$ in GPC can be identified as the so-called Q-value function in offline RL, as they predict the reward of an action chunk $a_{t:t+T}$ given the current information $I_{t}$.
+However, there are three subtle differences between GPC and offline RL.
+(a) GPC learns an explicit dynamics model instead of just approximating the Q-value function (this is similar to model-based offline RL).
+(b) The composition of ${\cal R}(\cdot)$ and ${\cal W}(\cdot)$ provides a Q-value function for “action chunks” while the usual Q-value function involves a single action.
+(c) GPC learns the dynamics model using diverse trajectories (without rewards), instead of only expert demonstrations.
+In §4, we compare GPC with an offline RL method adapted from single action to action chunk and show that GPC outperforms offline RL.
+In experiment section, we compare with an offline RL based planning baseline [31].
+While we rely on the world model to rank the action proposals, offline reinforcement learning ranks the action proposals using the value function [31].
+It firstly learns the value function via offline RL pre-training, and then performs test-time action re-ranking via the learned value function.
+To train a value function $Q(I_{t},a_{t:t+T})$ where $I_{t}$ is the state and $a_{t:t+T}$ is the action chunk, we need to obtain a reward function that can be used to supervise the value function training.
+Consistent with the GPC-RANK method, we use registration error between the object and the target at $I_{t+T+1}$ as the reward.
+We use the same length of action chunk ($T$) as GPC-RANK for a fair comparison.
+Implicit Q-learning (IQL) [41] is used as the offline RL method to pretrain a value function $Q(I_{t},a_{t:t+T})$.
+For vision-based tasks, we use the pretrained vision encoder (ResNet18) from the action policy to encode the observation images $I_{t}$ to latent vectors $s_{t}$.
+We then train the value function $Q(s_{t},a_{t:t+T})$.
+With the trained value function, we can rank and select the action proposal with the highest Q-value.
+$a_{t:t+T}^{(k\star)}=\operatorname*{arg\,max}_{a_{t:t+T}^{(k)},k=1\ldots K}Q(s_{t},a_{t:t+T}^{(k)}),$ (A3)
+which is the same as the GPC-RANK method in (2).
+Appendix F Real-World Experimental Results
+Here we provide more results for real-world experiments.
+F.1 Test Trajectories
+In this section, we provide examples of the test trajectories for five real-world tasks.
+The test trajectories for push-T with collision of A and R, clothes folding can be found in §4.3.
+Here we present trajectories for plain push-T, push-T with collision of A, and push-T in an unseen testing scenarios.
+More videos for all test seeds can be found in https://computationalrobotics.seas.harvard.edu/GPC/.
+(a) Test for push-T
+(b) Test for push-T with collision of A
+Figure A1: Real-world tests for plain push-T and push-T with collision.
+In every test, top row shows trajectories of baseline model ($K=1,M=0$), middle row shows trajectories of GPC-RANK ($K=10,M=0$), and last row shows trajectories of GPC-OPT ($K=0,M=25$).
+The first and last column show the initial and final position of the object, respectively.
+The middle five columns show overlaid trajectories generated by the policies.
+Figure A2: Real-world test for out-of-domain scene.
+In every test, top row shows trajectories of baseline model ($K=1,M=0$), middle row shows trajectories of GPC-RANK ($K=10,M=0$), and last row shows trajectories of GPC-OPT ($K=0,M=25$).
+This displays GPC would help to make policy succeed in unseen scene while baseline failed.
+F.2 Quality of world model predictions
+Fig A3, Fig A4, Fig A5, Fig A6, and Fig A7 show the quality of world model predictions in GPC.
 Refer to captionRefer to caption
@@
 (b) World model prediction in GPC-OPT
-Figure 6: World model prediction in GPC for push-T.
-[Top] For GPC-RANK, we show the action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$.
-Every row corresponds to the set of image consequences for an action proposal.
-We only pick $6$ action proposals and downsample the image prediction sequences for displaying purpose.
-[Bottom] For GPC-OPT, we show optimized action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$.
-Every row corresponds to the set of image consequences for an optimized action proposal at optimization step $i$.
-We only pick $6$ optimization steps and downsample the image prediction sequences for displaying purpose.
-Quality of world model predictions.
-Fig.
-6 shows the quality of world model prediction of push T for GPC-RANK and GPC-OPT.
-The images shown are all predicted by the world model.
-They not only have very good visual effect and look realistic, but also they can accurately predict the dynamics reaction from given actions.
-From the last column, we can see even when the differences between action sequences are very small, the world model can still capture the details and predict the reaction in a meticulous way.
-This is very important to make GPC work.
-IV-C2 Push-T under collision
-Next, we hope to investigate how GPC works under more complex scenarios, involving harder tasks and more complex dynamics.
-We put an object of shape “A” on top of the target position so that when pushing T, it will inevitably collide with object A.
-The training data is $100$ push-T expert demos with collision of A and some random exploration demos.
-Performance. We test $5$ seeds in total, and baseline model, GPC-RANK, GPC-OPT succeed for $2$, $4$, $3$ seeds respectively.
-We then proceed to test in an even harder situation: we put objects of shape “A” and “R” on top of the target position at the same time.
-Case 1: When pushing T, T may collide with A and R at the same time.
-Case 2: When pushing T, T would collide with one letter (e.g., “R”) and this letter will collide with another letter (e.g., “A”).
-We train the policy and world model here using $100$ plain push-T expert demonstrations, $100$ push-T colliding A expert demonstrations, and $100$ push-T colliding A&R expert demonstrations.
-Performance. In the total of $5$ testing seeds, baseline model, GPC-RANK, GPC-OPT succeed $2$, $3$, $4$ respectively.
-Fig. 10(b)(c) show the trajectories for both of these tests. More testing trajectories are shown in Appendix -B and supplementary videos.
-Refer to captionRefer to caption
-(a) World model prediction in GPC-RANK
-Refer to captionRefer to caption
-(b) World model prediction in GPC-OPT
-Figure 7: World model prediction in GPC for push-T with collision of A.
+Figure A3: World model prediction in GPC for push-T.
 [Top] For GPC-RANK, we show the action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$.
@@
 (b) World model prediction in GPC-OPT
-Figure 8: World model prediction in GPC for push-T with collision of A and R.
+Figure A4: World model prediction in GPC for push-T with collision of A.
 [Top] For GPC-RANK, we show the action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$.
@@
 We only pick $6$ optimization steps and downsample the image prediction sequences for displaying purpose.
-Quality of world model predictions.
-Fig.
-7 and Fig.
-8 show the quality of world model prediction for GPC-RANK and GPC-OPT, having collisions during push-T.
-The images shown are all predicted by the world model.
-With adding some collision objects, the world model is not only able to give near perfect prediction (i.e.
-no deformation) for a complex scene composing several objects, but also it can accurately predict the dynamics among these objects.
-IV-C3 Generalization to unseen cases
-Last but not least, we want to understand whether GPC can help when generalizing to unseen cases.
-We use the policy and world model trained with push-T, push-T colliding A, push-T colliding A&R expert demos.
-But we perform the test to push T when an object of letter “R” is on top of the target.
-This situation has never been seen in any of the training data.
-We test three different starting positions for T and R.
-Performance.
-GPC-RANK and GPC-OPT succeed in all of the three tests while baseline model fails in one of the cases.
-The reason behind this is that when generalizing to unseen cases, the policy’s performance degrades.
-So we need to either generate more samples from the action policy’s distribution (like giving more opportunities to generate an optimal proposal) or do gradient ascent to optimize the action proposal.
-Fig. 11 shows one test. More testing trajectories are shown in Appendix -B and supplementary videos.
 Refer to captionRefer to caption
@@
 (b) World model prediction in GPC-OPT
-Figure 9: World model prediction in GPC for push-T with collision of R.
+Figure A5: World model prediction in GPC for push-T with collision of A and R.
 [Top] For GPC-RANK, we show the action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$.
@@
 We only pick $6$ optimization steps and downsample the image prediction sequences for displaying purpose.
-Quality of world model predictions.
-Fig.
-9 shows the quality of world model prediction of push T under collision in unseen scenarios for GPC-RANK and GPC-OPT.
-A scenario only composes of T and R is never seen in the training data.
-However, our world model can still give good predictions for action sequences here.
-It verifies the ability of generalization for our world model.
-Remark 3 (Hallucination).
-One of the important issues of generative modeling is hallucination, and our visual world model is no exception.
-In particular, there are several cases where the world model predicts images of the future that violates rigid-body physics.
-For example, the objects T, A, and R may penetrate each other, as we show in Appendix -C.
-Interestingly, even though the predictions can violate physics, using such predictions for planning still leads to GPC’s superior performance.
-We believe learning a visual world model that strictly satisfies the laws of physics remains a challenging task, and strategies such as data scaling and baking in physics priors may help solve this challenge.
-(a) Test for push-T
-(b) Test for push-T with collision of A
-(c) Test for push-T with collision of A and R
-Figure 10: Real-world tests for plain push-T and push-T with collision.
-In every test, top row shows trajectories of baseline model ($K=1,M=0$), middle row shows trajectories of GPC-RANK ($K=10,M=0$), and last row shows trajectories of GPC-OPT ($K=0,M=25$).
-The first and last column show the initial and final position of the object, respectively.
-The middle five columns show overlaid trajectories generated by the policies.
-Results for other real-world tests are shown in Appendix -B and supplementary videos.
-Figure 11: Real-world test for out-of-domain scene.
-In every test, top row shows trajectories of baseline model ($K=1,M=0$), middle row shows trajectories of GPC-RANK ($K=10,M=0$), and last row shows trajectories of GPC-OPT ($K=0,M=25$).
-This displays GPC would help to make policy succeed in unseen scene while baseline failed.
-Remaining results for push-T with collision of R are shown in Appendix -B and supplementary videos.
-V Related Work
+Refer to captionRefer to caption
+(a) World model prediction in GPC-RANK
+Refer to captionRefer to caption
+(b) World model prediction in GPC-OPT
+Figure A6: World model prediction in GPC for push-T with collision of R.
+[Top] For GPC-RANK, we show the action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$.
+Every row corresponds to the set of image consequences for an action proposal.
+We only pick $6$ action proposals and downsample the image prediction sequences for displaying purpose.
+[Bottom] For GPC-OPT, we show optimized action proposals (left) and corresponding image predictions (right) from time $t$ to $t+T$.
+Every row corresponds to the set of image consequences for an optimized action proposal at optimization step $i$.
+We only pick $6$ optimization steps and downsample the image prediction sequences for displaying purpose.
+Figure A7: World model prediction in GPC for clothes folding.
+Every row corresponds to the set of image consequences for an action proposal.
+We only pick $6$ action proposals and downsample the image prediction sequences for displaying purpose.
+The one in red box is picked by policy.
+To note that only in the selected action proposal, the gripper is grasping the clothes according to the predictions (last image in row 4), while in row 1-3, the gripper, in a closed status, does not managed to grasp anything and in row 5-6, the gripper is still open and not grasping.
+Appendix G Related Work
 Generative modeling for robotics.
@@
 Different from these works, our work uses a video model as an action-conditioned dynamics model that enables us to build a generative control framework that online computes a sequence of actions to optimizes a task-specific reward function.
-VI Conclusion
-We presented generative predictive control (GPC), a learning control paradigm that combines a generative policy with a predictive world model during online planning.
-It is worth noting that GPC bears great similarity with the classical literature on predictive control in the sense that the dynamics model is first identified and then optimization is performed online to decide actions.
-The new insights offered by GPC, perhaps, are that (a) modern conditional diffusion models provide a powerful tool to learn controllable action-conditioned visual world models; and (b) a generative policy trained from behavior cloning provides good warmstarts for online optimization that seem promising to partially tackle the nonconvexity of online optimization.
-Through an in-depth investigation on planar pushing—both simulated and real—we demonstrated the superior robustness of GPC in both state-based and vision-based robotic manipulation.
-Limitations.
-Two issues are worth highlighting.
-The first issue is already mentioned in Remark 3 and concerns the problem of hallucinating future predictions that violate the laws of physics.
-To address this, it may be possible to leverage the so-called world foundation models [2]—trained on Internet-scale videos—and finetune them for specific robotic tasks.
-The second issue is associated with the computational complexity of using an expensive diffusion-based world model in online planning.
-It is worth noting that, although GPC dominates the baseline in terms of success rates, it is generally slower than the baseline due to predicting the future using the world model.
-A promising solution is to leverage consistency models [51] and one-step diffusion [61] for faster world model prediction.
-Acknowledgments
-This work was partially supported by the Harvard University Dean’s Competitive Fund for Promising Scholarship.
-The authors thank William Zhang for his help on hardware design and prototyping.
-We thank members of the Harvard Computational Robotics Group for their contributions in collecting demonstrations.
-Their collective support has been instrumental in the completion of this work.